### v2

# Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Cameras

Abstract
The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid motion estimation. This study investigates the tremendous potential of spike cameras (a type of ultra-high-speed, high-dynamic-range camera) in PIV. We propose a deep learning framework, Spike Imaging Velocimetry (SIV), designed specifically for highly turbulent and intricate flow fields. To aggregate motion features from the spike stream while minimizing information loss, we incorporate a Detail-Preserving Hierarchical Transform (DPHT) module. Additionally, we introduce a Graph Encoder (GE) to extract contextual features from highly complex fluid flows. Furthermore, we present a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), which provides labeled data for three challenging fluid dynamics scenarios. Our proposed method achieves superior performance compared to existing baseline methods on PSSD. The datasets and our implementation of SIV are open-sourced in the supplementary materials.
1 Introduction
The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research. This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraftKompenhans et al. (2000), optimizing vehicle designs to reduce air resistanceNakagawa et al. (2016), improving energy conversion efficiency in combustion processesChen et al. (2022), understanding the impact of blood flow dynamics on diseasesRaschi et al. (2012), and predicting the spread of water pollutionMoayeriKashani et al. (2016).
In PIV experiments or simulations, tracer particles(e.g., silver-coated particles) are introduced into the fluid to scatter optical signals. By estimating their displacements, PIV can infer the velocities of the fluid in which they are suspended.
Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetSCai et al. (2019b) and RAFT-PIVLagemann et al. (2021).
The success of optical flow networksDosovitskiy et al. (2015); Sun et al. (2018); Teed and Deng (2020) inspires the development of learning-based PIV algorithmsCai et al. (2019b, a); Lagemann et al. (2021). However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field. Although optical flow networks generally exhibit good generalization ability, the characteristics of fluid scenes cannot be ignored for further performance improvement. Turbulence is an irregular flow phenomenon occurring at high Reynolds numbers. When measuring the turbulent velocity field with PIV, this scenario is significantly different from macroscopic worlds, with the following key challenges:
(1) The fluid often moves rapidly within the field of view.
(2) Signals are only present in regions with tracer particles, and since the particle volume is small, downsampling may lead to signal loss;
(3) Unlike solids, the fluid is highly unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
These challenges often coexist in turbulence. Challenge (1) places high demands on experimental hardware, typically requiring ultra-high-speed cameras. Challenge (2) requires us to carefully address potential resolution loss in the network. Challenge (3) introduces problems to the extraction of context features. Based on these characteristics, we propose improvements in both hardware and algorithm aspects.
In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV. We argue that using spike cameras instead of CMOS sensors in fluid measurement tasks offers several advantages:
High temporal resolution (HTR). Spike cameras can achieve ultra-high temporal resolution (20000Hz), reducing the interval $\Delta t$ between the source and target time, and providing rich temporal information for accurate predictions. This allows the spike camera to cope well with Challenge (1);
High dynamic range (HDR). When particles accumulate in large numbers, local regions may become overly bright, exceeding the dynamic range of conventional cameras. Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range compared to CMOS sensors. This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated areas);
Reduced experimental complexity. Spike cameras can continuously capture images at high frame rates with exceptionally low bandwidth and eliminate the need for synchronization controllers to modulate the camera and laser.
In terms of algorithm, we introduce a specialized network, Spike Imaging Velocimetry (SIV) to estimate fluid displacement fields. We propose two modules, the Detail-Preserving Hierarchical Transform (DPHT) designed in consideration of Challenge (2), and the Graph Encoder (GE) mainly designed for Challenge (3). We also introduce a dataset called Particle Scenes with Spike and Displacement (PSSD). PSSD comprises labeled data addressing three common motion estimation scenarios: steady turbulence (Problem 1), high-speed flow (Problem 2), and high dynamic range scenarios (Problem 3). In the experimental section, we conduct comparative analyses of our method against both image-based and spike-based approaches. Our proposed SIV achieves state-of-the-art performance across all three sub-datasets. The effectiveness of the proposed modules is further demonstrated through ablation studies. The main contributions of this work are as follows:
1. Explore a feasible approach for applying spike cameras to fluid velocity measurement;
2. From the perspective of fluid dynamics, especially turbulent flows, and the characteristics of the PIV scenarios, design a suitable network structure SIV, including two new modules: DPHT and GE;
3. Propose a PIV dataset for recording continuous fluid motion with spike cameras, which includes four flow types and three challenging experimental scenarios.
2 Related Work
2.1 Spike-based Optical Flow Estimation
The spike cameraDong et al. (2021); Huang (2019) is a type of neuromorphic camera (NeurCam). Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina. Spike cameras mimic the encoding principles of the central fovea in the retina, providing continuous visual information with ultra-high temporal resolutions, high dynamic range(HDR), and significantly reduced energy consumption and latency, making them well-suited for high-speed motion scenarios. Their pixels independently accumulate photons and emit a spike when the accumulated intensity exceeds a certain threshold. This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images. For a detailed explanation of the working mechanism, refer to Dong et al. (2021). The application of spike cameras in optical flow estimation provides significant insights for this study. Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of Particle Image Velocimetry (PIV). The techniques employed in PIV for image registration and displacement estimation are essentially a variant of optical flow estimation. Spike-based optical flow estimation is still a relatively new field. Hu et al. (2022) introduced a specialized neural network architecture for spike-based optical flow estimation. This structure can adaptively eliminate motion blur in spike streams based on prior motion information. A framework based on Teed and Deng (2020) is proposed by Zhao et al. (2022) to handle continuous spike streams and is then extended by Zhao et al. (2024) to suppress the adverse effects of randomness in spike streams. Chen et al. (2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by Xia et al. (2024).
2.2 Image-based PIV Algorithms
Many traditional image-based PIV algorithmsLourenco and Krothapalli (1995); Nobach and Honkanen (2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation areas (IA). Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent. However, traditional methods depend on finely tuned parameters and complex optimization techniques. With advancements in optical flow networks, various architecturesDosovitskiy et al. (2015); Sun et al. (2018); Teed and Deng (2020) are developed to predict dense displacement fields of macroscopic objects with high accuracy. Early PIV networks like PIV-DCNNLee et al. (2017) laid the groundwork for these developments. PIV-NetSCai et al. (2019b) is capable of generating full-resolution displacement outputs. Cai et al. (2019a) adapt LiteFlowNetHui et al. (2018) to PIV. RAFT-PIVLagemann et al. (2021), derived from RAFTTeed and Deng (2020), has shown exceptional performance on multiple datasets. Additionally, unsupervised learning frameworks for PIVZhang and Piggott (2020); Zhang et al. (2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements. Other recent PIV networks include LIMAManickathan et al. (2023), CC-FCNGao et al. (2021), En-FlowNetCJi et al. (2024a), ARAFT-FlowNetHan and Wang (2023), OFVNetSJi et al. (2024b), and PIV-PWCNetZhang et al. (2023b). However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field, and most of them have not been open-sourced.
2.3 Event-based PIV Algorithms
The event camera(such as DVSLichtsteiner et al. (2008), DAVISMoeys et al. (2017), ATISPosch et al. (2010), CeleXHuang et al. (2017)) is another type of NeurCam, which is inspired by the peripheral retinaRebecq et al. (2019). Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes. When the brightness change of a pixel exceeds a preset threshold, the camera records this change event (brightness increase as a positive event, brightness decrease as a negative event) and outputs its timestamp, position, and change value as data. However, due to the lack of absolute intensity values, event cameras struggle to reconstruct complete images, particularly in static regions and with fine textures. The application of event cameras to PIV has already been explored by Willert and Klinner (2022) and Willert (2023). Similar to the spike camera, the event camera offers higher temporal resolution, high dynamic range, low latency, and reduced experimental complexity. However, event cameras are only sensitive to objects with obvious motion and have hardly any output signal in stationary regions. This can be a weakness compared to spike cameras, considering that not all areas of the flow field may be moving at high speed.
3 Preliminary
3.1 The Working Principle of the Spike Camera
Each photosensitive element in a spike camera generates spike streams sequentially over time. A ”1” in the stream indicates that a spike has occurred (i.e., the threshold has been reached), while a ”0” indicates an ongoing accumulation state. The Integral-and-Fire (IF) mechanism of spike cameras can be mathematically represented as follows:
$A(\mathbf{x},t)=\int_{0}^{t}\alpha\cdot I(\mathbf{x},\tau)\bmod\theta,$ (1)
Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator. The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represent the time length of the spike stream. Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and brightness representation.
3.2 Data Generation
Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3). Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground truth turbulent displacement fields when $\Delta t=21$ (about 1.05 ms) and $\Delta t=11$ (about 0.55 ms). Additionally, we provide corresponding pairs of images for comparative analysis for image-based methods. We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream. The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB)Li et al. (2008). JHTDB employs Direct Numerical Simulation (DNS) to simulate turbulent flow fields. DNS directly solves the Navier-Stokes equations without any model assumptions or simplifications, enabling precise capture of all scales in turbulent flow. This method resolves the fluid motion in both space and time with high-resolution grids, accurately capturing the complex and multi-scale nature of turbulence. Additional details about the dataset can be found in our supplementary materials.
4 Approaches
4.1 Overall Architecture
Problem definition. Let $\mathbf{S}\in\mathbb{B}^{H\times W\times T}$ represent the spike stream recorded by the spike camera. $\mathbb{B}$ is the binary domain. We aim to estimate the displacement $\mathbf{D}$ of particles in the fluid between time moments $t$ and $t+\Delta t$. At the pixel level, for any given pixel $\mathbf{p}=(x,y)$ at $t$, we need to find a unique correspondence formulated by:
$\mathcal{I}(\mathbf{p},t)\leftarrow\mathcal{I}(\mathbf{p},t+\Delta t).$ (2)
Here, $\mathcal{I}$ is the mapping of the spike stream $\mathbf{S}$ in the high-dimensional feature space, representing the scene information, i.e., $\mathcal{I}=\mathrm{Rep}(\mathbf{S})$.
Network architecture. The primary structure of the SIV network is shown in Figure 2. To represent the scenes at moments $t$ and $t+\Delta t$, we extract spike stream segments (source and target) near these two moments, denoted as $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$, and input them into the network. Firstly, through the proposed DPHT module, the sub-stream $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$ are embedded into the representations $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ by our proposed DPHT module. Motion features $\mathbf{F}_{s,m}$ and $\mathbf{F}_{t,m}$ are extracted from $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$. Through the proposed Graph Encoder, the context feature $\mathbf{F}_{c}$ is extracted from $\mathbf{R}_{t}$. Subsequently, the enhanced features $\mathbf{F}_{s,m}$, $\mathbf{F}_{t,m}$ and $\mathbf{F}_{c}$ undergo an iterative optimization process in a pyramid structure, ranging from coarse (1/16 downsampled branch) to fine (1/4 downsampled branch). In each branch, similar to DIPZheng et al. (2022), the Inverse Patchmatch Block is used to construct the correlation volume, and is aggregated through a GRU. Finally, the Local Search Block estimates the residual displacement field $\Delta\mathbf{D}$ by searching for local cost based on the current displacement field estimate, and cost aggregation is performed through another GRU.
Figure 2: The overall network structure of SIV includes motion feature extraction supported by DPHT and the motion encoder, context feature extraction supported by DPHT and the Graph Encoder, along with a multi-scale iterative optimization process.
Figure 3: (a) A demonstration of the Detail-Preserving Hierarchical Transform (DPHT) module. (b) A brief illustration of Graph Encoder (GE) module.
4.2 Detail-Preserving Hierarchical Transform
Zhao et al. (2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network. Low-level features mainly focus on local details, while high-level features aggregate these low-level features to express more complex meanings. However, our ablation experiments (see Table 2) show that directly using multi-level representations does not achieve optimal performance improvement for PIV. This is because, as mentioned in the Introduction, the reduction in resolution or scale of feature maps leads to a significant loss of detail information. Furthermore, these details may not be critical for macroscopic objects, but considering that the particles in PIV are small and the fluid motion is highly intricate and complex, even the loss of detail information could result in an increase in local prediction errors. Moreover, the sparse and dark regions without particles essentially carry little to no signal, much of the information in a small region may stem from the bright particles. Based on these considerations, we designed a 3D Detail-Preserving Downsampling unit in which a bunch of 3D convolutions and 3D max poolings are incorporated (as shown in Figure 3(a)) to reduce the progressive loss of important information across the multi-layer structure while enabling the network to aggregate features at multiple scales. We refer to the entire block in Figure 3(a) as the Detail-Preserving Hierarchical Transform (DPHT).
3D Detail-Preserving downsampling. As shown in Figure 3(a), suppose the input of level $l$ is a feature vector $\mathbf{x}^{(l)}$, and the processing of $\mathbf{x}^{(l)}$ at this level is as follows:
$\mathbf{x}^{(l)}_{\mathrm{out}}=\mathscr{D}^{(l)}(\mathbf{x}^{(l)})=\mathscr{F}_{\mathrm{2C}}(\mathscr{F}_{\mathrm{3C}}(\mathbf{x}^{(l)})+\mathscr{F}_{\mathrm{3MP}}(\mathbf{x}^{(l)})),$ (3)
Here, $\mathscr{F}_{\mathrm{3C}}$ represents a 5×3×3 3D convolution, $\mathscr{F}_{\mathrm{3MP}}$ is a combination of a 3D max pooling operation and a 5×1×1 3D convolution, and $\mathscr{F}_{\mathrm{2C}}$ refers to a sequential stack of 2D convolutions. The output of the previous level serves as the input to the next level, i.e., $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}_{\mathrm{out}}=\mathscr{D}^{(l)}(\mathbf{x}^{(l)})$.
Global temporal aggregation. In this step, the output feature vectors from each level are further refined by several convolutions (for detailed information, refer to the supplementary materials), and then concatenated together to form the final representation of the DPHT:
$\mathbf{R}=\mathscr{A}_{m}[\mathbf{x}^{(l=1)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=2)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=3)}_{\mathrm{out}}],$ (4)
$\oplus$ represents concatenation, and $\mathscr{A}_{m}$ denotes convolution. As shown in Figure 3, we use DPHT as a pre-representation method before motion feature extraction and context information extraction, obtaining $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ for source sub-stream and target sub-tream, respectively.
4.3 Graph Encoder
Turbulent motion is highly complex, and extracting context information from turbulence is challenging. Therefore, Yu et al. (2021) argued that context extractors are unnecessary for PIV and removed this structure from their network. However, we hold a different view: despite the spatially chaotic nature of turbulence, there still exist multi-scale vortex structures (referred to as quasi-order structures), making the extraction of context information meaningful. To achieve this, we introduce Graph Convolutional Unit (GCU) in the context extractor, referring to the entire context extractor as the Graph Encoder. Graph convolutions weight the relationships between nodes using an adjacency matrix, performing local aggregation. Each node not only focuses on its own features but also passes information from neighboring nodes (as defined by the adjacency matrix) to itself. We believe that this approach is particularly suited for capturing spatial or topological structures in the PIV data, especially when the quasi-order structures of the flow is highly complex topologically. This situation arises particularly in cases with extremely high flow speeds (meaning high Reynolds numbers). When the Reynolds number is sufficiently high, turbulence experiences a cascade process from large-scale vortices to small-scale vortices (Kolmogorov scale) Kolmogorov (1991). This cascade process is the transfer of energy from large vortices to small vortices in turbulence, with each vortex disturbing the surrounding flow field, further complicating the flow.
Graph projection. The first step is to map the preliminary context information $\mathbf{R}_{c}=\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the graph space. The projection operation decouples the position information from the original grid features, making the resulting low-dimensional node feature representations more compact and expressive. The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}$, contain appearance features related to the shape and regional information of the scene context. Given $\mathbf{R}_{c}$, we can use the projection function $\mathscr{P}_{R\to V}(\cdot)$ to assign features with similar representations to the same node, resulting $\mathbf{V}_{c}\in\mathbb{R}^{C\times K}$, where $C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes. Since $\mathscr{P}_{R\to V}(\cdot)$ we define is a linear combination of feature vectors within the grid space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{c})$, the resulting nodes can theoretically aggregate both long-range and short-range information across the entire feature map, which could be beneficial for turbulent scenes with multi-scale quasi-order structures. This is formulated by,
$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{c,j}).$ (5)
Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{c})\in\mathbb{R}^{(H\times W)\times K}$.
Graph convolution. After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the adjacency matrix $\mathbf{A}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors. We utilize the graph convolution $\mathscr{G}$ from Selvan et al. (2020) to further propagate information across the graph. Specifically, for a single graph convolution with its parameter $W_{g}$, the operation is defined as
$\mathbf{V_{c}^{{}^{\prime}}}=\mathscr{G}(\mathbf{V}_{c},\mathbf{A})=\sigma(\mathbf{A}\mathbf{V}_{c}^{T}W_{g}),$ (6)
where $\sigma(\cdot)$ is a non-linear activation function. Two serial graph convolutional networks (GCNs) are used to capture the relationships between nodes. The node features updated by the GCNs are given by
$\mathbf{V_{c}^{\mathrm{out}}}=\mathscr{G}_{2}[\mathscr{G}_{1}(\mathbf{V}_{c},\mathbf{A}),\mathbf{A}],$ (7)
Graph re-projection. Our graph re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}=[z_{ij}]$ as inputs, generating a 2D feature map $\widetilde{\mathbf{R}}_{c}$. Since inverting the projection weight matrix is not feasible, we follow the approach in Li and Gupta (2018); Luo et al. (2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$. This is given by,
$\widetilde{\mathbf{R}}_{c}=\mathscr{P}_{V\to R}(\mathbf{V_{c}^{\mathrm{out}}},\mathbf{Z})=\mathbf{Z}\mathbf{V_{c}^{\mathrm{out}}}^{T}.$ (8)
Thus, $\mathscr{P}_{V\to R}$ thus linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters. Note that even if two pixels are assigned to the same node, they will have different features after re-projection. Therefore, the GCU may preserve the spatial details of the signal. Finally, these projection results can be integrated into an existing residual extractor network $\mathscr{R}$ introduced by Teed and Deng (2020). Final context feature $\mathbf{F}_{c}$ given by,
$\mathbf{F}_{c}=\mathscr{R}(\mathbf{R}_{c}+\alpha\widetilde{\mathbf{R}}_{c}),$ (9)
where $\alpha$ denotes a learnable parameter that is initialized as 0 and gradually performs a weighted sum.
5 Experimental Results
In this section, we evaluate our method on the PSSD dataset and compare its performance with other traditional and learning-based methods. Additionally, we conduct a series of ablation experiments to assess the effectiveness of the proposed modules.
Figure 4: Visualization examples of different algorithms in Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3) are shown. For each problem, the red box in the top optical flow map and the yellow box in the bottom EPE error map indicate the same region of the flow field.
5.1 Experimental Details
SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU. We employ the Adam optimizer with a warm-up strategy during training. The model parameters for DIP are consistent with those specified in the original paperZheng et al. (2022). For each problem, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs. The number of iterative refinements is set to 10, and the batch size is 4.
Following Cai et al. (2019b) and Lagemann et al. (2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms. Additionally, given the absence of a dedicated outlier metric in PIV, we introduced Root Mean Squared Error (RMSE) as a supplementary metric due to its sensitivity to outliers.
Since SIV outputs a prediction sequence of length $N$ at each iteration, the loss function is defined as:
$loss=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1},w_{i}=\gamma^{N-i-1}.$ (10)
We set $\gamma=0.8$ during training.
Methods Channel Isotropic Mhd Mixing Average
$\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$
Problem 1 SIV (ours) 0.242 / 0.284 0.157 / 0.182 0.918 / 1.288 0.574 / 0.765 0.638 / 0.934 0.407 / 0.563 0.700 / 1.076 0.433 / 0.616 0.625 / 0.895 0.393 / 0.532
RAFT-PIV - Image 0.377 / 0.437 0.215 / 0.249 2.181 / 2.769 1.252 / 1.574 1.336 / 1.701 0.842 / 1.072 1.875 / 2.486 1.064 / 1.394 1.442 / 1.848 0.843 / 1.072
RAFT-PIV - Spike 0.347 / 0.401 0.208 / 0.240 1.223 / 1.570 0.755 / 0.942 0.857 / 1.136 0.557 / 0.731 0.958 / 1.288 0.581 / 0.755 0.846 / 1.099 0.525 / 0.667
PIV-NetS - Image 0.545 / 0.904 0.301 / 0.478 2.560 / 3.136 1.497 / 1.819 1.574 / 2.121 0.985 / 1.284 2.247 / 2.845 1.320 / 1.651 1.731 / 2.253 1.026 / 1.308
PIV-NetS - Spike 0.553 / 0.821 0.289 / 0.424 1.507 / 1.991 0.874 / 1.124 1.252 / 1.782 0.734 / 1.019 1.382 / 1.903 0.782 / 1.040 1.173 / 1.624 0.670 / 0.902
SCFlow 0.354 / 0.468 0.222 / 0.284 2.232 / 2.706 1.305 / 1.628 1.374 / 1.703 0.901 / 1.110 2.132 / 2.684 1.169/ 1.435 1.522 / 1.890 0.899 / 1.114
HiST-SFlow 0.369 / 0.424 0.224 / 0.256 1.271 / 1.614 0.796 / 0.980 0.964 / 1.250 0.611 / 0.784 1.029 / 1.354 0.637 / 0.809 0.908 / 1.160 0.567 / 0.708
coarse-to-fine HS 6.790 / 8.448 0.509 / 0.655 7.415 / 9.354 1.998 / 3.079 7.683 / 10.783 2.558 / 4.463 9.388 / 7.274 1.954 / 3.146 7.819 / 9.751 1.755 / 2.836
Problem 2 SIV (ours) 0.366 / 0.443 0.256 / 0.300 1.721 / 2.602 1.076 / 1.521 1.098 / 1.671 0.717 / 1.026 1.482 / 2.493 0.907 / 1.412 1.167 /1.802 0.739 / 1.065
RAFT-PIV - Image 0.665 / 0.785 0.368 / 0.434 4.730 / 5.952 2.635 / 3.297 2.320 / 2.906 1.366 / 1.707 4.305 / 5.572 2.368 / 3.065 3.005 / 3.804 1.684 / 2.126
RAFT-PIV - Spike 0.688 / 0.792 0.412 / 0.477 2.410 / 3.169 1.483 / 1.882 1.468 / 1.936 0.959 / 1.236 2.071 / 2.864 1.241 / 1.649 1.659 / 2.190 1.024 / 1.311
PIV-NetS - Image 1.163 / 1.896 0.683 / 1.029 4.866 / 5.944 2.830 / 3.409 2.673 / 3.719 1.643 / 2.207 4.347 / 5.497 2.500 / 3.095 3.262 / 4.264 1.914 / 2.435
PIV-NetS - Spike 0.887 / 1.552 0.540 / 0.836 3.172 / 4.155 1.818 / 2.315 2.392 / 3.400 1.410 / 1.963 3.067 / 4.154 1.744 / 2.316 2.379 / 3.315 1.378 / 1.857
SCFlow 0.694 / 0.716 0.403 / 0.486 4.689 / 5.975 2.613 / 3.156 2.387 / 3.055 1.391 / 1.697 4.326 / 5.262 2.380 / 3.123 3.024 / 3.752 1.696 / 2.116
HiST-SFlow 0.556 / 0.641 0.364 / 0.419 2.212 / 2.947 1.365 / 1.756 1.404 / 1.879 0.998 / 1.308 1.911 / 2.693 1.179 / 1.595 1.521 / 2.040 0.976 / 1.269
coarse-to-fine HS 20.691 / 20.993 8.244 / 9.628 17.164 / 19.627 8.051 / 9.944 17.469 / 21.594 8.170 / 11.252 17.131 / 19.572 7.936 / 9.970 18.114 / 20.447 8.100 / 10.199
Problem 3 SIV (ours) 0.241 / 0.286 0.155 / 0.181 0.903 / 1.282 0.564 / 0.761 0.621 / 0.916 0.393 / 0.542 0.698 / 1.106 0.437 / 0.634 0.616 / 0.897 0.389 / 0.530
RAFT-PIV - Image 0.447 / 0.523 0.244 / 0.286 2.694 / 3.334 1.495 / 1.847 1.447 / 1.808 0.888 / 1.104 2.311 / 2.948 1.271 / 1.613 1.725 / 2.153 0.974 / 1.213
RAFT-PIV - Spike 0.320 / 0.373 0.200 / 0.232 1.203 / 1.558 0.740 / 0.939 0.798 / 1.061 0.530 / 0.698 0.946 / 1.297 0.576 / 0.770 0.817 / 1.072 0.512 / 0.660
PIV-NetS - Image 0.667 / 0.991 0.343 / 0.513 2.786 / 3.357 1.602 / 1.924 1.641 / 2.164 0.956 / 1.258 2.393 / 2.988 1.378 / 1.709 1.872 / 2.375 1.070 / 1.351
PIV-NetS - Spike 0.476 / 0.789 0.247 / 0.406 1.420 / 1.914 0.856 / 1.134 1.157 / 1.695 0.716 / 1.021 1.342 / 1.891 0.774 / 1.063 1.099 / 1.572 0.648 / 0.906
SCFlow 0.462 / 0.743 0.269 / 0.433 1.850 / 2.206 0.837 / 1.008 1.300 / 1.915 0.774 / 1.142 1.972 / 2.618 0.769 / 1.040 1.396 / 1.871 0.662 / 0.905
HiST-SFlow 0.329 / 0.381 0.180 / 0.210 1.179 / 1.523 0.640 / 0.828 0.832 / 1.093 0.452 / 0.616 0.960 / 1.302 0.476 / 0.663 0.825 / 1.075 0.437 / 0.579
coarse-to-fine HS 7.908 / 8.946 1.044 / 1.648 7.924 / 9.634 2.706 / 3.816 8.345 / 11.002 3.130 / 4.828 7.854 / 9.668 2.716 / 3.903 8.008 / 9.813 2.399 / 3.549
[…]
5.2 Comparative Experiments
To evaluate the effectiveness of our proposed method, we conduct extensive experiments on the PSSD dataset. The experiments are performed under identical settings to ensure fairness in the results.
Index Settings Problem 1 Problem 2 Problem 3
DPHT GE $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$
(A) 0.625 0.393 1.167 0.739 0.616 0.389
(B) ✓* 0.658 0.395 1.249 0.764 0.627 0.389
(C) 0.757 0.403 1.278 0.777 0.630 0.391
(D) ✓* 0.645 0.396 1.242 0.752 0.619 0.386
(E) 0.701 0.398 1.249 0.793 0.659 0.401
(F) 0.805 0.431 1.349 0.887 0.725 0.413
Table 2: Ablation study of our proposed modules, showing results for End-Point Error (EPE). Blocks colored greener indicate better performance, characterized by lower EPE values. ✓ indicates the presence of the module, ✗ indicates the absence, and ✓* indicates the removal of the submodules under this module, for DPHT, it refers to 3D-DPD, and for GE, it refers to GCU (including the projection and re-projection). Setting (A) represents our final model.
We compare our method with multiple open-source baselines, including a traditional PIV algorithmCai et al. (2018), PIV networks (PIV-NetSCai et al. (2019b), RAFT-PIVLagemann et al. (2021)), as well as their spike versions, and spike-based optical flow networks (SCFlowHu et al. (2022), HiST-SFlowZhao et al. (2024)).
Results on Problem 1. In the steady flow scenario, the fluid moves at a moderate speed. In both the time intervals of 1 and 2, SIV achieves the smallest EPE and RMSE, providing the optimal estimation. Compared to the suboptimal methods, SIV reduces the EPE by 26.1% when $\Delta t=21$. The traditional coarse-to-fine HSCai et al. (2018) method can only yield relatively reasonable predictions when the time interval and displacement are small, which reflects, to some extent, the challenge of the dataset. Figure 4 shows the displacement field predictions (optical flow maps) and error maps for different methods in PIV. The red boxes in the optical flow map correspond to the yellow boxes in the error map. The marked region in Problem 1 is a smooth boundary between the two components. By comparing with the ground truth optical flow map, it is clear that the optical flow map shapes predicted by SIV and the spike version of RAFT-PIVLagemann et al. (2021) are mostly accurate, while other methods show varying degrees of distortion in this region. In the error map, it is more intuitive to see that the red areas (high error regions) in SIV are much fewer.
Results on Problem 2. Problem 2 represents high-speed scenarios with greater particle displacements. SIV maintains its lead in all test scenarios, with a 23.3% reduction in EPE at $\Delta t=21$ compared to the suboptimal method. The traditional method struggles to make effective predictions in this scenario. In the visualization example of Problem 2 shown in Figure 4, the marked area includes a fine flow field shaped like a ”human hand” and a ”laminar flow region” with very high flow speed. The predicted flow field shape of SIV is very close to the ground truth, and the error in the high-speed ”laminar flow region” is significantly smaller than that of other methods.
Results on Problem 3. Problem 3 reflects scenarios with uneven lighting, where areas may exist that cause overexposure in traditional cameras (this issue does not occur in spike cameras due to their principle). SIV still achieves the best performance across all sub-datasets. In the visualization example of Problem 3 (see Figure 4), only SIV predicts a flow field in the marked region that is nearly consistent with the ground truth, and has the fewest red areas (high error regions) in the error map. As we anticipated, in Problem 3, spike-based methods outperform image-based methods, reflecting the high dynamic range advantage of spike cameras under uneven lighting conditions.
5.3 Ablation Study
We conduct a series of ablation experiments to evaluate the contribution of each component to overall performance, as detailed in Table 2. Each column in the table is color-coded with a gradient from yellow to green, where the greener color represents better performance. These results underscore the effectiveness of the DPHT and Graph Encoder (GE) and their submodules. In Problem 1, both DPHT and GE exhibit notable improvements. In high-speed motion scenarios (Problem 2), particles in the fluid exhibit greater displacements between the source time and target time, increasing the difficulty of accurate flow prediction. Under these circumstances, the enhancements provided by the 3D-DPD submodule in DPHT and the GCU submodule in GE (7.0% and 6.4%, respectively) are the most significant across the three Problems. In Problem 3, the distinctive brightness distribution of the scene might aid feature extraction, and the improvements from DPHT and GE are comparatively less pronounced but still remain observable.
6 Conclusions
We start from the perspective of fluid (primarily turbulent) scenarios and conduct a detailed analysis of their prominent characteristics, which may pose challenges to PIV hardware and deep learning algorithms. Based on these analyses, we explore a novel approach to estimate fluid motion using spike cameras. Additionally, we develop a tailored network architecture called SIV, designed to solve PIV problems while adapting to scenario-specific features, which integrates two novel modules: DPHT and GE. DPHT is mainly used to aggregate motion features from the spike stream while retaining scene details. GE is primarily used to extract contextual features from highly unstructured turbulent scenes. We also propose a more challenging and realistic PIV dataset, PSSD, which features multiple scenarios and captures the temporal dynamics of flow field variations. Through comparative experiments, we validate that our method outperforms existing open-source PIV algorithms and spike-based optical flow algorithms on the PSSD dataset. In ablation studies, we gradually remove the introduced modules and their submodules to prove their effectiveness and analyze their performance in different scenarios. In summary, this study explores the great potential of spike cameras in PIV and advocates for the use of customized network solutions for fluid motion estimation tasks.

### v3

@@
-Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Cameras
+Spike Imaging Velocimetry: Dense Motion Estimation of Fluids Using Spike Streams
 Abstract
-The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid motion estimation.
-This study investigates the tremendous potential of spike cameras (a type of ultra-high-speed, high-dynamic-range camera) in PIV.
-We propose a deep learning framework, Spike Imaging Velocimetry (SIV), designed specifically for highly turbulent and intricate flow fields.
-To aggregate motion features from the spike stream while minimizing information loss, we incorporate a Detail-Preserving Hierarchical Transform (DPHT) module.
-Additionally, we introduce a Graph Encoder (GE) to extract contextual features from highly complex fluid flows.
-Furthermore, we present a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), which provides labeled data for three challenging fluid dynamics scenarios.
-Our proposed method achieves superior performance compared to existing baseline methods on PSSD.
-The datasets and our implementation of SIV are open-sourced in the supplementary materials.
-1 Introduction
-The need for accurate and non-intrusive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research.
-This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraftKompenhans et al.
-(2000), optimizing vehicle designs to reduce air resistanceNakagawa et al.
-(2016), improving energy conversion efficiency in combustion processesChen et al.
-(2022), understanding the impact of blood flow dynamics on diseasesRaschi et al.
-(2012), and predicting the spread of water pollutionMoayeriKashani et al.
-(2016).
+Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows.
+It is commonly employed to analyze complex flow structures and validate numerical simulations.
+This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry.
+We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation.
+To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR).
+Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics.
+Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness.
+The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
+Abstract
+Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows.
+It is commonly employed to analyze complex flow structures and validate numerical simulations.
+This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry.
+We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation.
+To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR).
+Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics.
+Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness.
+The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
+Introduction
+The need for accurate and non-invasive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research.
+This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraft (Kompenhans et al.
+2000), optimizing vehicle designs to reduce air resistance (Nakagawa et al.
+2016), improving energy conversion efficiency in combustion processes (Chen, Li, and Zhu 2022), understanding the impact of blood flow dynamics on diseases (Raschi et al.
+2012), and predicting the spread of water pollution (MoayeriKashani et al.
+2016).
+Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetS (Cai et al.
+2019b) and RAFT-PIV (Lagemann et al.
+2021).
 In PIV experiments or simulations, tracer particles (e.g., silver-coated particles) are introduced into the fluid to scatter optical signals.
 By estimating their displacements, PIV can infer the velocities of the fluid in which they are suspended.
-Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetSCai et al.
-(2019b) and RAFT-PIVLagemann et al.
-(2021).
-The success of optical flow networksDosovitskiy et al.
-(2015); Sun et al.
-(2018); Teed and Deng (2020) inspires the development of learning-based PIV algorithmsCai et al.
-(2019b, a); Lagemann et al.
-(2021).
-However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field.
+The success of optical flow networks (Dosovitskiy et al.
+2015; Sun et al.
+2018; Teed and Deng 2020) has inspired the development of learning-based PIV algorithms (Cai et al.
+2019b, a; Lagemann et al.
+2021).
 Although optical flow networks generally exhibit good generalization ability, the characteristics of fluid scenes cannot be ignored for further performance improvement.
@@
 (1) The fluid often moves rapidly within the field of view.
-(2) Signals are only present in regions with tracer particles, and since the particle volume is small, downsampling may lead to signal loss;
-(3) Unlike solids, the fluid is highly unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
+(2) The scene contains extensive empty regions without tracer particles, yet the network is required to estimate the full flow field.
+(3) Unlike solids, the fluid is highly sparse and unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
 These challenges often coexist in turbulence.
-Challenge (1) places high demands on experimental hardware, typically requiring ultra-high-speed cameras.
-Challenge (2) requires us to carefully address potential resolution loss in the network.
+Challenge (1) places high demands on the experimental hardware, typically requiring ultra-high-speed cameras.
+Challenge (2) requires addressing potential resolution loss in the network carefully.
 Challenge (3) introduces problems to the extraction of context features.
-Based on these characteristics, we propose improvements in both hardware and algorithm aspects.
-In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV.
-We argue that using spike cameras instead of CMOS sensors in fluid measurement tasks offers several advantages:
+To address these challenges, we propose improvements in both hardware and algorithm aspects.
+In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV. We argue that using spike cameras in fluid measurement tasks offers several advantages:
 High temporal resolution (HTR).
 Spike cameras can achieve ultra-high temporal resolution (20000Hz), reducing the interval $\Delta t$ between the source and target time, and providing rich temporal information for accurate predictions.
-This allows the spike camera to cope well with Challenge (1);
+This enables spike cameras to handle Challenge (1) effectively;
 High dynamic range (HDR).
 When particles accumulate in large numbers, local regions may become overly bright, exceeding the dynamic range of conventional cameras.
-Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range compared to CMOS sensors.
-This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated areas);
+Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range.
+This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated regions);
 Reduced experimental complexity.
 Spike cameras can continuously capture images at high frame rates with exceptionally low bandwidth and eliminate the need for synchronization controllers to modulate the camera and laser.
-In terms of algorithm, we introduce a specialized network, Spike Imaging Velocimetry (SIV) to estimate fluid displacement fields.
-We propose two modules, the Detail-Preserving Hierarchical Transform (DPHT) designed in consideration of Challenge (2), and the Graph Encoder (GE) mainly designed for Challenge (3).
-We also introduce a dataset called Particle Scenes with Spike and Displacement (PSSD).
-PSSD comprises labeled data addressing three common motion estimation scenarios: steady turbulence (Problem 1), high-speed flow (Problem 2), and high dynamic range scenarios (Problem 3).
-In the experimental section, we conduct comparative analyses of our method against both image-based and spike-based approaches.
-Our proposed SIV achieves state-of-the-art performance across all three sub-datasets.
-The effectiveness of the proposed modules is further demonstrated through ablation studies.
-The main contributions of this work are as follows:
-1. Explore a feasible approach for applying spike cameras to fluid velocity measurement;
-2. From the perspective of fluid dynamics, especially turbulent flows, and the characteristics of the PIV scenarios, design a suitable network structure SIV, including two new modules: DPHT and GE;
-3. Propose a PIV dataset for recording continuous fluid motion with spike cameras, which includes four flow types and three challenging experimental scenarios.
-2 Related Work
-2.1 Spike-based Optical Flow Estimation
-The spike cameraDong et al.
-(2021); Huang (2019) is a type of neuromorphic camera (NeurCam).
+We propose a specialized network, Spike Imaging Velocimetry (SIV), for estimating fluid velocity fields.
+To enhance network performance, we develop three novel modules tailored to the characteristics of fluid dynamics and spike cameras.
+These include the Detail-Preserving Hierarchical Transform (DPHT), which enhances feature extraction from spike streams while mitigating downsampling artifacts; the Graph Encoder (GE), which maps features into a graph structure for adaptive context aggregation; and the Multi-scale Velocity Refinement (MSVR), which addresses over-smoothing in predicted flow fields.
+We also introduce a new dataset, Particle Scenes with Spike and Displacement (PSSD), designed for learning fluid motion from spike data.
+It covers three representative motion estimation scenarios: steady turbulence, high-speed flow, and high dynamic range conditions.
+Extensive experiments demonstrate the superiority of SIV over both image-based and spike-based baselines.
+SIV achieves state-of-the-art performance across all evaluation settings.
+The effectiveness of each proposed module is further validated through ablation studies.
+The main contributions of this work are summarized as follows:
+1.
+We explore a feasible solution for applying spike cameras to fluid velocity measurement, opening new directions for PIV.
+2. We design a dedicated network, SIV, tailored to fluid dynamics and turbulent flow properties, and propose three novel modules: the DPHT, GE, and MSVR.
+3. We propose a new dataset, PSSD, which captures continuous fluid motion across four flow types and three challenging experimental scenarios.
+4.
+We conduct extensive quantitative experiments and qualitative visual comparisons, demonstrating that SIV achieves state-of-the-art performance across multiple benchmarks.
+We further perform ablation studies to systematically validate the effectiveness of each proposed module.
+Related Work
+Spike-based Optical Flow Estimation
+The spike camera (Dong, Huang, and Tian 2021; Huang 2019) is a type of neuromorphic camera (NeurCam).
 Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina.
@@
 This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images.
-For a detailed explanation of the working mechanism, refer to Dong et al.
-(2021).
+For a detailed explanation of the working mechanism, refer to (Dong, Huang, and Tian 2021).
 The application of spike cameras in optical flow estimation provides significant insights for this study.
-Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of Particle Image Velocimetry (PIV).
+Although optical flow estimation is generally used to estimate the motion of macroscopic objects, its principles and objectives are aligned with those of PIV.
 The techniques employed in PIV for image registration and displacement estimation are essentially a variant of optical flow estimation.
 Spike-based optical flow estimation is still a relatively new field.
-Hu et al.
-(2022) introduced a specialized neural network architecture for spike-based optical flow estimation.
+(Hu et al.
+2022) introduced a specialized neural network architecture for spike-based optical flow estimation.
 This structure can adaptively eliminate motion blur in spike streams based on prior motion information.
-A framework based on Teed and Deng (2020) is proposed by Zhao et al.
-(2022) to handle continuous spike streams and is then extended by Zhao et al.
-(2024) to suppress the adverse effects of randomness in spike streams.
-Chen et al.
-(2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by Xia et al.
-(2024).
-2.2 Image-based PIV Algorithms
-Many traditional image-based PIV algorithmsLourenco and Krothapalli (1995); Nobach and Honkanen (2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation areas (IA).
+A framework based on RAFT is proposed by (Zhao et al.
+2022) to handle continuous spike streams.
+(Zhao et al.
+2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network.
+(Chen, Yu, and Huang 2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by (Xia et al.
+2024).
+Image-based PIV Algorithms
+Many traditional image-based PIV algorithms (Lourenco and Krothapalli 1995; Nobach and Honkanen 2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation windows.
 Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent.
 However, traditional methods depend on finely tuned parameters and complex optimization techniques.
-With advancements in optical flow networks, various architecturesDosovitskiy et al.
-(2015); Sun et al.
-(2018); Teed and Deng (2020) are developed to predict dense displacement fields of macroscopic objects with high accuracy.
-Early PIV networks like PIV-DCNNLee et al.
-(2017) laid the groundwork for these developments.
-PIV-NetSCai et al.
-(2019b) is capable of generating full-resolution displacement outputs.
-Cai et al.
-(2019a) adapt LiteFlowNetHui et al.
-(2018) to PIV.
-RAFT-PIVLagemann et al.
-(2021), derived from RAFTTeed and Deng (2020), has shown exceptional performance on multiple datasets.
-Additionally, unsupervised learning frameworks for PIVZhang and Piggott (2020); Zhang et al.
-(2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements.
-Other recent PIV networks include LIMAManickathan et al.
-(2023), CC-FCNGao et al.
-(2021), En-FlowNetCJi et al.
-(2024a), ARAFT-FlowNetHan and Wang (2023), OFVNetSJi et al.
-(2024b), and PIV-PWCNetZhang et al.
-(2023b).
-However, the performance improvements of these PIV networks mainly rely on advancements in optical flow estimation networks, rather than targeted improvements that incorporate the specific characteristics of the flow field, and most of them have not been open-sourced.
-2.3 Event-based PIV Algorithms
-The event camera(such as DVSLichtsteiner et al.
-(2008), DAVISMoeys et al.
-(2017), ATISPosch et al.
-(2010), CeleXHuang et al.
-(2017)) is another type of NeurCam, which is inspired by the peripheral retinaRebecq et al.
-(2019).
+With advancements in optical flow networks, various architectures (Dosovitskiy et al.
+2015; Sun et al.
+2018; Teed and Deng 2020) have been developed to predict dense velocity fields of macroscopic objects with high accuracy.
+Early PIV networks like PIV-DCNN (Lee, Yang, and Yin 2017) laid the groundwork for these developments.
+PIV-NetS (Cai et al.
+2019b) is capable of generating full-resolution displacement outputs.
+(Cai et al.
+2019a) adapt LiteFlowNet (Hui, Tang, and Loy 2018) to PIV.
+RAFT-PIV (Lagemann et al.
+2021), derived from RAFT (Teed and Deng 2020), has shown exceptional performance on multiple datasets.
+(Yu et al.
+2021) removes the computationally inefficient components of RAFT, such as the context encoder, to improve inference speed.
+Additionally, unsupervised learning frameworks for PIV(Zhang and Piggott 2020; Zhang et al.
+2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements.
+Other recent PIV networks include LIMA (Manickathan, Mucignat, and Lunati 2023), CC-FCN (Gao et al.
+2021), En-FlowNetC (Ji, An, and Hui 2024), ARAFT-FlowNet (Han and Wang 2023), OFVNetS (Ji, Hui, and An 2024), and PIV-PWCNet (Zhang et al.
+2023b).
+However, the performance gains of most PIV networks are largely attributed to advances in optical flow estimation architectures, rather than targeted enhancements tailored to the specific characteristics of fluid dynamics, and most implementations remain unavailable as open-source.
+Event-based PIV Algorithms
+The event camera (Lichtsteiner, Posch, and Delbruck 2008; Moeys et al.
+2017; Posch, Matolin, and Wohlgenannt 2010; Huang, Guo, and Chen 2017) is another type of NeurCam, which is inspired by the peripheral retina (Rebecq et al.
+2019).
 Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes.
@@
 However, due to the lack of absolute intensity values, event cameras struggle to reconstruct complete images, particularly in static regions and with fine textures.
-The application of event cameras to PIV has already been explored by Willert and Klinner (2022) and Willert (2023).
+The application of event cameras to PIV has already been explored by (Willert and Klinner 2022) and (Willert 2023).
 Similar to the spike camera, the event camera offers higher temporal resolution, high dynamic range, low latency, and reduced experimental complexity.
-However, event cameras are only sensitive to objects with obvious motion and have hardly any output signal in stationary regions.
-This can be a weakness compared to spike cameras, considering that not all areas of the flow field may be moving at high speed.
-3 Preliminary
-3.1 The Working Principle of the Spike Camera
+Preliminary
+The Working Principle of the Spike Camera
 Each photosensitive element in a spike camera generates spike streams sequentially over time.
@@
 Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator.
-The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represent the time length of the spike stream.
+The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represents the temporal dimension.
 Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and brightness representation.
-3.2 Data Generation
+Data Generation
 Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3).
-Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground truth turbulent displacement fields when $\Delta t=21$ (about 1.05 ms) and $\Delta t=11$ (about 0.55 ms).
+Each sub-dataset comprises 10,000 sample sequences of spike streams and the corresponding ground-truth turbulent velocity fields when $\Delta t=21$ (approximately 1.05 ms) and $\Delta t=11$ (approximately 0.55 ms).
 Additionally, we provide corresponding pairs of images for comparative analysis for image-based methods.
 We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream.
-The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB)Li et al.
-(2008).
+The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB) (Li et al.
+2008).
 JHTDB employs Direct Numerical Simulation (DNS) to simulate turbulent flow fields.
@@
 Additional details about the dataset can be found in our supplementary materials.
-4 Approaches
-4.1 Overall Architecture
+Approaches
+Overall Architecture
 Problem definition.
@@
 $\mathcal{I}(\mathbf{p},t)\leftarrow\mathcal{I}(\mathbf{p},t+\Delta t).$ (2)
-Here, $\mathcal{I}$ is the mapping of the spike stream $\mathbf{S}$ in the high-dimensional feature space, representing the scene information, i.e., $\mathcal{I}=\mathrm{Rep}(\mathbf{S})$.
-Network architecture.
-The primary structure of the SIV network is shown in Figure 2.
-To represent the scenes at moments $t$ and $t+\Delta t$, we extract spike stream segments (source and target) near these two moments, denoted as $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$, and input them into the network.
-Firstly, through the proposed DPHT module, the sub-stream $\mathbf{S}_{s}$ and $\mathbf{S}_{t}$ are embedded into the representations $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ by our proposed DPHT module.
-Motion features $\mathbf{F}_{s,m}$ and $\mathbf{F}_{t,m}$ are extracted from $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$.
-Through the proposed Graph Encoder, the context feature $\mathbf{F}_{c}$ is extracted from $\mathbf{R}_{t}$.
-Subsequently, the enhanced features $\mathbf{F}_{s,m}$, $\mathbf{F}_{t,m}$ and $\mathbf{F}_{c}$ undergo an iterative optimization process in a pyramid structure, ranging from coarse (1/16 downsampled branch) to fine (1/4 downsampled branch).
-In each branch, similar to DIPZheng et al.
-(2022), the Inverse Patchmatch Block is used to construct the correlation volume, and is aggregated through a GRU.
-Finally, the Local Search Block estimates the residual displacement field $\Delta\mathbf{D}$ by searching for local cost based on the current displacement field estimate, and cost aggregation is performed through another GRU.
-Figure 2: The overall network structure of SIV includes motion feature extraction supported by DPHT and the motion encoder, context feature extraction supported by DPHT and the Graph Encoder, along with a multi-scale iterative optimization process.
-Figure 3: (a) A demonstration of the Detail-Preserving Hierarchical Transform (DPHT) module. (b) A brief illustration of Graph Encoder (GE) module.
-4.2 Detail-Preserving Hierarchical Transform
-Zhao et al.
-(2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network.
-Low-level features mainly focus on local details, while high-level features aggregate these low-level features to express more complex meanings.
-However, our ablation experiments (see Table 2) show that directly using multi-level representations does not achieve optimal performance improvement for PIV.
-This is because, as mentioned in the Introduction, the reduction in resolution or scale of feature maps leads to a significant loss of detail information.
-Furthermore, these details may not be critical for macroscopic objects, but considering that the particles in PIV are small and the fluid motion is highly intricate and complex, even the loss of detail information could result in an increase in local prediction errors.
-Moreover, the sparse and dark regions without particles essentially carry little to no signal, much of the information in a small region may stem from the bright particles.
-Based on these considerations, we designed a 3D Detail-Preserving Downsampling unit in which a bunch of 3D convolutions and 3D max poolings are incorporated (as shown in Figure 3(a)) to reduce the progressive loss of important information across the multi-layer structure while enabling the network to aggregate features at multiple scales.
-We refer to the entire block in Figure 3(a) as the Detail-Preserving Hierarchical Transform (DPHT).
+Here, $\mathcal{I}$ is the mapping of the spike stream $\mathbf{S}$ in the high-dimensional feature space, representing the scene information.
+Figure 2: Spike Imaging Velocimetry (SIV) network.
+Figure 3: (a) Detail-Preserving Hierarchical Transform (DPHT) Module. (b) Graph Encoder (GE).
+Detail-Preserving Hierarchical Transform
+In order to avoid the loss of particle signal information during feature extraction, we propose a multi-scale pyramidal architecture, named Detail-Preserving Hierarchical Transform (DPHT), which incorporates a 3D Detail-Preserving Downsampling module (as shown in Figure 3(a)).
 3D Detail-Preserving downsampling. As shown in Figure 3(a), suppose the input of level $l$ is a feature vector $\mathbf{x}^{(l)}$, and the processing of $\mathbf{x}^{(l)}$ at this level is as follows:
@@
 In this step, the output feature vectors from each level are further refined by several convolutions (for detailed information, refer to the supplementary materials), and then concatenated together to form the final representation of the DPHT:
-$\mathbf{R}=\mathscr{A}_{m}[\mathbf{x}^{(l=1)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=2)}_{\mathrm{out}}\oplus\mathbf{x}^{(l=3)}_{\mathrm{out}}],$ (4)
-$\oplus$ represents concatenation, and $\mathscr{A}_{m}$ denotes convolution.
-As shown in Figure 3, we use DPHT as a pre-representation method before motion feature extraction and context information extraction, obtaining $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$ for source sub-stream and target sub-tream, respectively.
-4.3 Graph Encoder
-Turbulent motion is highly complex, and extracting context information from turbulence is challenging.
-Therefore, Yu et al.
-(2021) argued that context extractors are unnecessary for PIV and removed this structure from their network.
-However, we hold a different view: despite the spatially chaotic nature of turbulence, there still exist multi-scale vortex structures (referred to as quasi-order structures), making the extraction of context information meaningful.
-To achieve this, we introduce Graph Convolutional Unit (GCU) in the context extractor, referring to the entire context extractor as the Graph Encoder.
-Graph convolutions weight the relationships between nodes using an adjacency matrix, performing local aggregation.
-Each node not only focuses on its own features but also passes information from neighboring nodes (as defined by the adjacency matrix) to itself.
-We believe that this approach is particularly suited for capturing spatial or topological structures in the PIV data, especially when the quasi-order structures of the flow is highly complex topologically.
-This situation arises particularly in cases with extremely high flow speeds (meaning high Reynolds numbers).
-When the Reynolds number is sufficiently high, turbulence experiences a cascade process from large-scale vortices to small-scale vortices (Kolmogorov scale) Kolmogorov (1991).
-This cascade process is the transfer of energy from large vortices to small vortices in turbulence, with each vortex disturbing the surrounding flow field, further complicating the flow.
+$\mathbf{R}=\mathscr{A}_{m}[\mathbf{Concat}(\mathbf{x}^{(l=1)}_{\mathrm{out}},\mathbf{x}^{(l=2)}_{\mathrm{out}},\mathbf{x}^{(l=3)}_{\mathrm{out}})$], (4)
+$\mathscr{A}_{m}$ denotes convolution. After processing with DPHT, the signals of the source and target sub-streams are denoted as $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$, respectively.
+Graph Encoder
+When the Reynolds number is high, fluid flow becomes turbulent.
+As energy cascades and dissipates, large vortices become unstable and gradually break into smaller ones until they vanish.
+This indicates that fluid motion is not purely chaotic, but rather exhibits an underlying topological structure.
+To capture the contextual information of the fluid, we propose a Graph Encoder (GE).
+It is a feature extraction network based on Graph Neural Networks (GNNs) and Convolutional Neural Networks (CNNs).
+By representing features as graph nodes and constructing a dynamic adjacency matrix, GE flexibly aggregates and updates information from neighboring nodes, enabling a larger receptive field than conventional CNNs.
+Graph Attention Networks (GATs) (Veličković et al.
+2017), leveraging attention mechanisms, learn the relative contribution of each neighbor to a node’s new representation, thereby mitigating interference from particle-sparse regions.
+The serial combination of GNN and CNN enables adaptive global perception and local refinement, effectively capturing fluid-specific contextual features.
 Graph projection.
-The first step is to map the preliminary context information $\mathbf{R}_{c}=\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the graph space.
-The projection operation decouples the position information from the original grid features, making the resulting low-dimensional node feature representations more compact and expressive.
-The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}$, contain appearance features related to the shape and regional information of the scene context.
-Given $\mathbf{R}_{c}$, we can use the projection function $\mathscr{P}_{R\to V}(\cdot)$ to assign features with similar representations to the same node, resulting $\mathbf{V}_{c}\in\mathbb{R}^{C\times K}$, where $C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes.
-Since $\mathscr{P}_{R\to V}(\cdot)$ we define is a linear combination of feature vectors within the grid space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{c})$, the resulting nodes can theoretically aggregate both long-range and short-range information across the entire feature map, which could be beneficial for turbulent scenes with multi-scale quasi-order structures.
-This is formulated by,
-$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{c,j}).$ (5)
-Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{c})\in\mathbb{R}^{(H\times W)\times K}$.
-Graph convolution.
-After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the adjacency matrix $\mathbf{A}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors.
-We utilize the graph convolution $\mathscr{G}$ from Selvan et al.
-(2020) to further propagate information across the graph.
-Specifically, for a single graph convolution with its parameter $W_{g}$, the operation is defined as
-$\mathbf{V_{c}^{{}^{\prime}}}=\mathscr{G}(\mathbf{V}_{c},\mathbf{A})=\sigma(\mathbf{A}\mathbf{V}_{c}^{T}W_{g}),$ (6)
-where $\sigma(\cdot)$ is a non-linear activation function.
-Two serial graph convolutional networks (GCNs) are used to capture the relationships between nodes.
-The node features updated by the GCNs are given by
-$\mathbf{V_{c}^{\mathrm{out}}}=\mathscr{G}_{2}[\mathscr{G}_{1}(\mathbf{V}_{c},\mathbf{A}),\mathbf{A}],$ (7)
+The first step is to map $\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the node space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{s})$.
+The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}\in\mathbb{R}^{C\times K}$, contain appearance features related to the shape and regional information of the scene.
+$C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes.
+In this paper, we set $K=128$.
+$\mathscr{P}_{R\to V}(\cdot)$ is a linear combination of feature vectors within the grid space, formulated by,
+$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{s,j}).$ (5)
+Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{s})\in\mathbb{R}^{(H\times W)\times K}$.
+Graph convolution layer.
+After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the dynamic adjacency matrix $\widetilde{\mathbf{A}}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors.
+The detailed definition of GAT can be found in the supplementary material.
+Here, we implicitly denote the operations of the two GAT modules as $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$.
+The updated node features are given by
+$\mathbf{V_{c}^{\mathrm{out}}}=\mathscr{G}_{2}[\mathscr{G}_{1}(\mathbf{V}_{c},\widetilde{\mathbf{A}})+\mathbf{V}_{c},\widetilde{\mathbf{A}}]+\mathscr{G}_{1}(\mathbf{V}_{c},\widetilde{\mathbf{A}})+\mathbf{V}_{c},$ (6)
 Graph re-projection.
 Our graph re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}=[z_{ij}]$ as inputs, generating a 2D feature map $\widetilde{\mathbf{R}}_{c}$.
-Since inverting the projection weight matrix is not feasible, we follow the approach in Li and Gupta (2018); Luo et al.
-(2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$.
+Since inverting the projection weight matrix is not feasible, we follow the approach in (Li and Gupta 2018; Luo et al.
+2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$.
 This is given by,
 $\widetilde{\mathbf{R}}_{c}=\mathscr{P}_{V\to R}(\mathbf{V_{c}^{\mathrm{out}}},\mathbf{Z})=\mathbf{Z}\mathbf{V_{c}^{\mathrm{out}}}^{T}.$ (7)
-Thus, $\mathscr{P}_{V\to R}$ thus linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters.
+Thus, $\mathscr{P}_{V\to R}$ linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters.
 Note that even if two pixels are assigned to the same node, they will have different features after re-projection.
 Therefore, the GCU may preserve the spatial details of the signal.
-Finally, these projection results can be integrated into an existing residual extractor network $\mathscr{R}$ introduced by Teed and Deng (2020).
+Finally, these projection results can be integrated into a residual extractor network $\mathscr{R}$ for local refinement.
 Final context feature $\mathbf{F}_{c}$ given by,
-$\mathbf{F}_{c}=\mathscr{R}(\mathbf{R}_{c}+\alpha\widetilde{\mathbf{R}}_{c}),$ (9)
+$\mathbf{F}_{c}=\mathscr{R}(\mathbf{R}_{s}+\alpha\widetilde{\mathbf{R}}_{c}),$ (8)
 where $\alpha$ denotes a learnable parameter that is initialized as 0 and gradually performs a weighted sum.
-5 Experimental Results
+Multi-scale Iterative Optimizer
+We adopt the Multi-scale Iterative Optimizer (MSIO) from DIP (Zheng et al.
+2022) as the backbone for iterative residual flow estimation in SIV.
+Unlike RAFT’s full correlation volume construction, MSIO leverages Inverse PatchMatch to significantly reduce the computational cost introduced by large-scale feature maps (1/4 resolution of the full size).
+The detailed architecture of MSIO can be found in (Zheng et al.
+2022); here, we denote it implicitly as $\mathscr{U}$.
+The output multi-scale velocity field is given by
+$\{\mathbf{Vel}^{i}\}_{i=1}^{\text{iters}}=\mathscr{U}(\mathbf{F}_{c},\mathbf{F}_{m}),$ (9)
+$\mathbf{F}_{m}$ denotes the motion features extracted by the Motion Encoder.
+iter denotes the number of iterations of GRU (Teed and Deng 2020).
+The detailed architecture of the Motion Encoder is provided in the supplementary material.
+Multi-scale Velocity Refinement
+DIP adopts the same compact and effective convex upsampling strategy as RAFT to recover full-resolution optical flow.
+However, such upsampling is insufficient for accurately reconstructing the velocity field of fluids.
+To address this, we propose a Multi-scale Velocity Refinement (MSVR) module to assist in reconstructing the complete velocity field, including fine-scale vortices.
+As illustrated in Figure 4, we first apply a sequence of 2D convolutions followed by activation functions to process the multi-scale velocity fields generated by MCIO.
+Information exchange across scales is achieved via cross convolution from coarse to fine levels.
+The extracted features are then fused and processed by distinct convolutional networks to generate a residual velocity field $\mathbf{Vel}^{\text{res}}$ and a quality map $\mathbf{Q}$.
+The refined fluid velocity field is given by,
+$\mathbf{Vel}^{*}=\mathbf{Vel}^{i=\text{iters}}+\mathbf{Vel}^{\text{res}}\odot\mathbf{Q},$ (10)
+$\odot$ denotes element-wise multiplication.
+Figure 4: Multi-scale Velocity Refinement Module.
+Loss Function
+Since SIV outputs a prediction sequence of length $N$ at each iteration, the flow loss is defined as:
+$L_{\text{flow}}=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1}.$ (11)
+We introduce a gradient loss to prevent the generated flow field from becoming overly smooth in regions like shear layers or small vortical structures. It is computed as follows:
+$L_{\text{grad}}=\sum_{i=0}^{i=N}w_{i}\cdot(\|\nabla_{\mathbf{x}}f_{i}-\nabla_{\mathbf{x}}f_{gt}\|_{1}+\|\nabla_{\mathbf{y}}f_{i}-\nabla_{\mathbf{y}}f_{gt}\|_{1}).$ (12)
+We set $w_{i}=\gamma^{N-i-1}$ and $\gamma=0.8$ for both $L_{\text{flow}}$ and $L_{\text{grad}}$. The total loss can be formulated as follows:
+$L=L_{\text{flow}}+\beta\cdot L_{\text{grad}}.$ (13)
+Here, we set $\beta=0.3$.
+Experimental Results
 In this section, we evaluate our method on the PSSD dataset and compare its performance with other traditional and learning-based methods.
@@
 Figure 5: Visualization examples of different algorithms in Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3) are shown.
-For each problem, the red box in the top optical flow map and the yellow box in the bottom EPE error map indicate the same region of the flow field.
-5.1 Experimental Details
+For each problem, the red box in the top optical flow map and the green box in the bottom EPE error map indicate the same region of the flow field.
+Experimental Details
 SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU.
 We employ the Adam optimizer with a warm-up strategy during training.
-The model parameters for DIP are consistent with those specified in the original paperZheng et al.
-(2022).
-For each problem, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs.
-The number of iterative refinements is set to 10, and the batch size is 4.
-Following Cai et al.
-(2019b) and Lagemann et al.
-(2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms.
-Additionally, given the absence of a dedicated outlier metric in PIV, we introduced Root Mean Squared Error (RMSE) as a supplementary metric due to its sensitivity to outliers.
-Since SIV outputs a prediction sequence of length $N$ at each iteration, the loss function is defined as:
-$loss=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1},w_{i}=\gamma^{N-i-1}.$ (10)
-We set $\gamma=0.8$ during training.
+For each sub-dataset, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs.
+The number of iterative refinements of GRUs is set to 12.
+The batch size is set to 4.
+Following (Cai et al.
+2019b) and (Lagemann et al.
+2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms.
+Additionally, we incorporate another widely used PIV metric, Root Mean Squared Error (RMSE), in our extensive experiments, as detailed in the supplementary material.
 Methods Channel Isotropic Mhd Mixing Average
 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11
-Problem 1 SIV (ours) 0.242 / 0.284 0.157 / 0.182 0.918 / 1.288 0.574 / 0.765 0.638 / 0.934 0.407 / 0.563 0.700 / 1.076 0.433 / 0.616 0.625 / 0.895 0.393 / 0.532
-RAFT-PIV - Image 0.377 / 0.437 0.215 / 0.249 2.181 / 2.769 1.252 / 1.574 1.336 / 1.701 0.842 / 1.072 1.875 / 2.486 1.064 / 1.394 1.442 / 1.848 0.843 / 1.072
-RAFT-PIV - Spike 0.347 / 0.401 0.208 / 0.240 1.223 / 1.570 0.755 / 0.942 0.857 / 1.136 0.557 / 0.731 0.958 / 1.288 0.581 / 0.755 0.846 / 1.099 0.525 / 0.667
-PIV-NetS - Image 0.545 / 0.904 0.301 / 0.478 2.560 / 3.136 1.497 / 1.819 1.574 / 2.121 0.985 / 1.284 2.247 / 2.845 1.320 / 1.651 1.731 / 2.253 1.026 / 1.308
-PIV-NetS - Spike 0.553 / 0.821 0.289 / 0.424 1.507 / 1.991 0.874 / 1.124 1.252 / 1.782 0.734 / 1.019 1.382 / 1.903 0.782 / 1.040 1.173 / 1.624 0.670 / 0.902
-SCFlow 0.354 / 0.468 0.222 / 0.284 2.232 / 2.706 1.305 / 1.628 1.374 / 1.703 0.901 / 1.110 2.132 / 2.684 1.169/ 1.435 1.522 / 1.890 0.899 / 1.114
-HiST-SFlow 0.369 / 0.424 0.224 / 0.256 1.271 / 1.614 0.796 / 0.980 0.964 / 1.250 0.611 / 0.784 1.029 / 1.354 0.637 / 0.809 0.908 / 1.160 0.567 / 0.708
-coarse-to-fine HS 6.790 / 8.448 0.509 / 0.655 7.415 / 9.354 1.998 / 3.079 7.683 / 10.783 2.558 / 4.463 9.388 / 7.274 1.954 / 3.146 7.819 / 9.751 1.755 / 2.836
-Problem 2 SIV (ours) 0.366 / 0.443 0.256 / 0.300 1.721 / 2.602 1.076 / 1.521 1.098 / 1.671 0.717 / 1.026 1.482 / 2.493 0.907 / 1.412 1.167 /1.802 0.739 / 1.065
-RAFT-PIV - Image 0.665 / 0.785 0.368 / 0.434 4.730 / 5.952 2.635 / 3.297 2.320 / 2.906 1.366 / 1.707 4.305 / 5.572 2.368 / 3.065 3.005 / 3.804 1.684 / 2.126
-RAFT-PIV - Spike 0.688 / 0.792 0.412 / 0.477 2.410 / 3.169 1.483 / 1.882 1.468 / 1.936 0.959 / 1.236 2.071 / 2.864 1.241 / 1.649 1.659 / 2.190 1.024 / 1.311
-PIV-NetS - Image 1.163 / 1.896 0.683 / 1.029 4.866 / 5.944 2.830 / 3.409 2.673 / 3.719 1.643 / 2.207 4.347 / 5.497 2.500 / 3.095 3.262 / 4.264 1.914 / 2.435
-PIV-NetS - Spike 0.887 / 1.552 0.540 / 0.836 3.172 / 4.155 1.818 / 2.315 2.392 / 3.400 1.410 / 1.963 3.067 / 4.154 1.744 / 2.316 2.379 / 3.315 1.378 / 1.857
-SCFlow 0.694 / 0.716 0.403 / 0.486 4.689 / 5.975 2.613 / 3.156 2.387 / 3.055 1.391 / 1.697 4.326 / 5.262 2.380 / 3.123 3.024 / 3.752 1.696 / 2.116
-HiST-SFlow 0.556 / 0.641 0.364 / 0.419 2.212 / 2.947 1.365 / 1.756 1.404 / 1.879 0.998 / 1.308 1.911 / 2.693 1.179 / 1.595 1.521 / 2.040 0.976 / 1.269
-coarse-to-fine HS 20.691 / 20.993 8.244 / 9.628 17.164 / 19.627 8.051 / 9.944 17.469 / 21.594 8.170 / 11.252 17.131 / 19.572 7.936 / 9.970 18.114 / 20.447 8.100 / 10.199
-Problem 3 SIV (ours) 0.241 / 0.286 0.155 / 0.181 0.903 / 1.282 0.564 / 0.761 0.621 / 0.916 0.393 / 0.542 0.698 / 1.106 0.437 / 0.634 0.616 / 0.897 0.389 / 0.530
-RAFT-PIV - Image 0.447 / 0.523 0.244 / 0.286 2.694 / 3.334 1.495 / 1.847 1.447 / 1.808 0.888 / 1.104 2.311 / 2.948 1.271 / 1.613 1.725 / 2.153 0.974 / 1.213
-RAFT-PIV - Spike 0.320 / 0.373 0.200 / 0.232 1.203 / 1.558 0.740 / 0.939 0.798 / 1.061 0.530 / 0.698 0.946 / 1.297 0.576 / 0.770 0.817 / 1.072 0.512 / 0.660
-PIV-NetS - Image 0.667 / 0.991 0.343 / 0.513 2.786 / 3.357 1.602 / 1.924 1.641 / 2.164 0.956 / 1.258 2.393 / 2.988 1.378 / 1.709 1.872 / 2.375 1.070 / 1.351
-PIV-NetS - Spike 0.476 / 0.789 0.247 / 0.406 1.420 / 1.914 0.856 / 1.134 1.157 / 1.695 0.716 / 1.021 1.342 / 1.891 0.774 / 1.063 1.099 / 1.572 0.648 / 0.906
-SCFlow 0.462 / 0.743 0.269 / 0.433 1.850 / 2.206 0.837 / 1.008 1.300 / 1.915 0.774 / 1.142 1.972 / 2.618 0.769 / 1.040 1.396 / 1.871 0.662 / 0.905
-HiST-SFlow 0.329 / 0.381 0.180 / 0.210 1.179 / 1.523 0.640 / 0.828 0.832 / 1.093 0.452 / 0.616 0.960 / 1.302 0.476 / 0.663 0.825 / 1.075 0.437 / 0.579
-coarse-to-fine HS 7.908 / 8.946 1.044 / 1.648 7.924 / 9.634 2.706 / 3.816 8.345 / 11.002 3.130 / 4.828 7.854 / 9.668 2.716 / 3.903 8.008 / 9.813 2.399 / 3.549
-[…]
-5.2 Comparative Experiments
-To evaluate the effectiveness of our proposed method, we conduct extensive experiments on the PSSD dataset. The experiments are performed under identical settings to ensure fairness in the results.
-Index Settings Problem 1 Problem 2 Problem 3
-DPHT GE $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$ $\Delta t=21$ $\Delta t=11$
-(A) 0.625 0.393 1.167 0.739 0.616 0.389
-(B) ✓* 0.658 0.395 1.249 0.764 0.627 0.389
-(C) 0.757 0.403 1.278 0.777 0.630 0.391
-(D) ✓* 0.645 0.396 1.242 0.752 0.619 0.386
-(E) 0.701 0.398 1.249 0.793 0.659 0.401
-(F) 0.805 0.431 1.349 0.887 0.725 0.413
-Table 2: Ablation study of our proposed modules, showing results for End-Point Error (EPE).
-Blocks colored greener indicate better performance, characterized by lower EPE values.
-✓ indicates the presence of the module, ✗ indicates the absence, and ✓* indicates the removal of the submodules under this module, for DPHT, it refers to 3D-DPD, and for GE, it refers to GCU (including the projection and re-projection).
-Setting (A) represents our final model.
-We compare our method with multiple open-source baselines, including a traditional PIV algorithmCai et al.
-(2018), PIV networks (PIV-NetSCai et al.
-(2019b), RAFT-PIVLagemann et al.
-(2021)), as well as their spike versions, and spike-based optical flow networks (SCFlowHu et al.
-(2022), HiST-SFlowZhao et al.
-(2024)).
-Results on Problem 1.
-In the steady flow scenario, the fluid moves at a moderate speed.
-In both the time intervals of 1 and 2, SIV achieves the smallest EPE and RMSE, providing the optimal estimation.
-Compared to the suboptimal methods, SIV reduces the EPE by 26.1% when $\Delta t=21$.
-The traditional coarse-to-fine HSCai et al.
-(2018) method can only yield relatively reasonable predictions when the time interval and displacement are small, which reflects, to some extent, the challenge of the dataset.
-Figure 4 shows the displacement field predictions (optical flow maps) and error maps for different methods in PIV.
-The red boxes in the optical flow map correspond to the yellow boxes in the error map.
-The marked region in Problem 1 is a smooth boundary between the two components.
-By comparing with the ground truth optical flow map, it is clear that the optical flow map shapes predicted by SIV and the spike version of RAFT-PIVLagemann et al.
-(2021) are mostly accurate, while other methods show varying degrees of distortion in this region.
-In the error map, it is more intuitive to see that the red areas (high error regions) in SIV are much fewer.
-Results on Problem 2.
-Problem 2 represents high-speed scenarios with greater particle displacements.
-SIV maintains its lead in all test scenarios, with a 23.3% reduction in EPE at $\Delta t=21$ compared to the suboptimal method.
-The traditional method struggles to make effective predictions in this scenario.
-In the visualization example of Problem 2 shown in Figure 4, the marked area includes a fine flow field shaped like a ”human hand” and a ”laminar flow region” with very high flow speed.
-The predicted flow field shape of SIV is very close to the ground truth, and the error in the high-speed ”laminar flow region” is significantly smaller than that of other methods.
-Results on Problem 3.
-Problem 3 reflects scenarios with uneven lighting, where areas may exist that cause overexposure in traditional cameras (this issue does not occur in spike cameras due to their principle).
-SIV still achieves the best performance across all sub-datasets.
-In the visualization example of Problem 3 (see Figure 4), only SIV predicts a flow field in the marked region that is nearly consistent with the ground truth, and has the fewest red areas (high error regions) in the error map.
-As we anticipated, in Problem 3, spike-based methods outperform image-based methods, reflecting the high dynamic range advantage of spike cameras under uneven lighting conditions.
-5.3 Ablation Study
+coarse-to-fine HS 6.790 0.509 7.415 1.998 7.683 2.558 9.388 1.954 7.819 1.755
+RAFT-PIV-Image 0.377 0.215 2.181 1.252 1.336 0.842 1.875 1.064 1.442 0.843
+RAFT-PIV-Spike 0.347 0.208 1.223 0.755 0.857 0.557 0.958 0.581 0.846 0.525
+PIV-NetS-Image 0.545 0.301 2.560 1.497 1.574 0.985 2.247 1.320 1.731 1.026
+PIV-NetS-Spike 0.553 0.289 1.507 0.874 1.252 0.734 1.382 0.782 1.173 0.670
+SCFlow 0.354 0.222 2.232 1.305 1.374 0.901 2.132 1.169 1.522 0.899
+HiST-SFlow 0.369 0.224 1.271 0.796 0.964 0.611 1.029 0.637 0.908 0.567
+Flowformer-Spike 0.286 0.147 1.069 0.642 0.727 0.437 0.806 0.472 0.722 0.425
+SIV (ours) 0.239 0.159 0.896 0.587 0.613 0.422 0.680 0.441 0.607 0.402
+Table 1: EPE ($\downarrow$) results on PSSD dataset (Problem 1). The best results are marked in bold.
+Comparative Experiments
+We compare our method against several open-source baselines, including a classical PIV algorithm (Cai et al.
+2018), PIV networks (PIV-NetS (Cai et al.
+2019b), RAFT-PIV (Lagemann et al.
+2021)), spike-adapted general optical flow networks (e.g., Flowformer with spike input (Huang et al.
+2022)), and spike-based optical flow networks (SCFlow (Hu et al.
+2022), HiST-SFlow (Zhao et al.
+2024)).
+Methods Channel Isotropic Mhd Mixing Average
+$\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11
+coarse-to-fine HS 20.691 8.244 17.164 8.051 17.469 8.170 17.131 7.936 18.114 8.100
+RAFT-PIV-Image 0.665 0.368 4.730 2.635 2.320 1.366 4.305 2.368 3.005 1.684
+RAFT-PIV-Spike 0.688 0.412 2.410 1.483 1.468 0.959 2.071 1.241 1.659 1.024
+PIV-NetS-Image 1.163 0.683 4.866 2.830 2.673 1.643 4.347 2.500 3.262 1.914
+PIV-NetS-Spike 0.887 0.540 3.172 1.818 2.392 1.410 3.067 1.744 2.379 1.378
+SCFlow 0.694 0.403 4.689 2.613 2.387 1.391 4.326 2.380 3.024 1.696
+HiST-SFlow 0.556 0.364 2.212 1.365 1.404 0.998 1.911 1.179 1.521 0.976
+Flowformer-Spike 0.533 0.284 1.975 1.179 1.214 0.748 1.668 0.972 1.348 0.796
+SIV (ours) 0.370 0.254 1.687 1.080 1.071 0.722 1.449 0.912 1.144 0.742
+Table 2: EPE ($\downarrow$) results on PSSD dataset (Problem 2). The best results are marked in bold.
+As shown in Table 1, 2, and 3, $\Delta t=21$ and $\Delta t=11$ represent the number of timestamps for the central temporal intervals between two spike streams used for velocity field prediction, respectively.
+For detailed descriptions, please refer to the supplementary material.
+Our proposed SIV method achieves the lowest average End-Point Error (EPE) across all evaluated sub-datasets.
+Compared to baseline approaches, SIV demonstrates superior prediction accuracy in both low-velocity (Problem 1) and high-velocity (Problem 2) optical flow scenarios.
+Furthermore, in high dynamic range conditions (Problem 3), SIV effectively exploits the inherent advantages of spike-based input modalities to address the fundamental limitations of conventional frame-based imaging systems.
+Figure 5 presents the visualized predictions of all methods across three datasets.
+The top row shows the velocity fields visualized using the standard optical flow color coding (Baker et al.
+2011).
+The bottom row shows the endpoint error (EPE) maps between the predicted and ground-truth flow.
+Shear layers are often a primary source of turbulence generation and development.
+SIV achieves the lowest prediction error in both shear layers and small-scale vortices, while avoiding noticeable blurring, distortion, or discontinuities.
+Methods Channel Isotropic Mhd Mixing Average
+$\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11
+coarse-to-fine HS 7.908 1.044 7.924 2.706 8.345 3.130 7.854 2.716 8.008 2.399
+RAFT-PIV-Image 0.447 0.244 2.694 1.495 1.447 0.888 2.311 1.271 1.725 0.974
+RAFT-PIV-Spike 0.320 0.200 1.203 0.740 0.798 0.530 0.946 0.576 0.817 0.512
+PIV-NetS-Image 0.667 0.343 2.786 1.602 1.641 0.956 2.393 1.378 1.872 1.070
+PIV-NetS-Spike 0.476 0.247 1.420 0.856 1.157 0.716 1.342 0.774 1.099 0.648
+SCFlow 0.462 0.269 1.850 0.837 1.300 0.774 1.972 0.769 1.396 0.662
+HiST-SFlow 0.329 0.180 1.179 0.640 0.832 0.452 0.960 0.476 0.825 0.437
+Flowformer-Spike 0.269 0.145 1.052 0.620 0.716 0.429 0.813 0.481 0.713 0.419
+SIV (ours) 0.234 0.155 0.880 0.573 0.598 0.404 0.677 0.438 0.598 0.392
+Table 3: EPE ($\downarrow$) results on PSSD dataset (Problem 3). The best results are marked in bold.
+Ablation Study
+Ablation of modules.
 We conduct a series of ablation experiments to evaluate the contribution of each component to overall performance, as detailed in Table 4.
-Each column in the table is color-coded with a gradient from yellow to green, where the greener color represents better performance.
-These results underscore the effectiveness of the DPHT and Graph Encoder (GE) and their submodules.
-In Problem 1, both DPHT and GE exhibit notable improvements.
-In high-speed motion scenarios (Problem 2), particles in the fluid exhibit greater displacements between the source time and target time, increasing the difficulty of accurate flow prediction.
-Under these circumstances, the enhancements provided by the 3D-DPD submodule in DPHT and the GCU submodule in GE (7.0% and 6.4%, respectively) are the most significant across the three Problems.
-In Problem 3, the distinctive brightness distribution of the scene might aid feature extraction, and the improvements from DPHT and GE are comparatively less pronounced but still remain observable.
-6 Conclusions
-We start from the perspective of fluid (primarily turbulent) scenarios and conduct a detailed analysis of their prominent characteristics, which may pose challenges to PIV hardware and deep learning algorithms.
-Based on these analyses, we explore a novel approach to estimate fluid motion using spike cameras.
-Additionally, we develop a tailored network architecture called SIV, designed to solve PIV problems while adapting to scenario-specific features, which integrates two novel modules: DPHT and GE.
-DPHT is mainly used to aggregate motion features from the spike stream while retaining scene details.
-GE is primarily used to extract contextual features from highly unstructured turbulent scenes.
-We also propose a more challenging and realistic PIV dataset, PSSD, which features multiple scenarios and captures the temporal dynamics of flow field variations.
-Through comparative experiments, we validate that our method outperforms existing open-source PIV algorithms and spike-based optical flow algorithms on the PSSD dataset.
-In ablation studies, we gradually remove the introduced modules and their submodules to prove their effectiveness and analyze their performance in different scenarios.
-In summary, this study explores the great potential of spike cameras in PIV and advocates for the use of customized network solutions for fluid motion estimation tasks.
+All the values are the arithmetic mean of the scenes.
+The best results are marked in bold.
+These results underscore the effectiveness of the DPHT, GE, MSVR and $L_{\text{grad}}$.
+In Table 4, the comparison between experiments (A) and (B) demonstrates the effectiveness of DPHT.
+Both the comparison between (B) and (C) and that between (D) and (E) demonstrate the effectiveness of the Graph Encoder (GE).
+Both the comparison between (B) and (D) and that between (C) and (E) highlight the contribution of MSVR.
+Ablation of $\beta$.
+Table 5 summarizes the ablation study on the weighting coefficient $\beta$ of the proposed loss function.
+The results indicate that setting $\beta=0.3$ yields the best performance for the current model and dataset.
+Additional ablation studies are provided in the supplementary material.
+# Settings Problem 1 Problem 2 Problem 3
+DPHT GE MSVR $\Delta t$=21 $\Delta t$=21 $\Delta t$=21
+(A) $\times$ $\times$ $\times$ 0.805 1.450 0.776
+(B) $\checkmark$ $\times$ $\times$ 0.701 1.299 0.659
+(C) $\checkmark$ $\checkmark$ $\times$ 0.625 1.167 0.616
+(D) $\checkmark$ $\times$ $\checkmark$ 0.644 1.268 0.635
+(E) $\checkmark$ $\checkmark$ $\checkmark$ 0.607 1.144 0.598
+Table 4: Ablation study of proposed modules, showing results for EPE ($\downarrow$). $\checkmark$ indicates module presence, $\times$ indicates absence. Setting (G) is our final model.
+# $\beta$ Problem 1 Problem 2 Problem 3
+(A) 0.0 0.620 1.169 0.615
+(B) 0.1 0.617 1.163 0.605
+(C) 0.2 0.615 1.157 0.603
+(D) 0.3 0.607 1.144 0.598
+(E) 0.4 0.613 1.146 0.610
+(F) 0.5 0.619 1.141 0.617
+Table 5: Ablation study of $\beta$, showing results for EPE (↓).
+Conclusion
+We explore a novel approach to fluid motion estimation using spike cameras and propose a specialized network architecture, termed Spike Imaging Velocimetry (SIV), designed to solve PIV tasks while adapting to the scenario-specific characteristics of fluid scenes.
+We propose a Detail-Preserving Hierarchical Transform module to preserve particle signal information during the pre-processing of spike streams.
+We design a Graph Encoder to extract contextual features from highly unstructured and turbulent flow fields.
+We develop a Multi-scale Velocity Refinement module to refine the velocity field.
+We also introduce a challenging and realistic PIV dataset, named Particle Scenes with Spike and Displacement (PSSD).
+Experimental results demonstrate that our method achieves state-of-the-art performance on the PSSD dataset and accurately estimates velocities in small-scale yet critical flow structures such as shear layers and vortices.
+In summary, this study demonstrates the potential of spike cameras in fluid velocity measurement and offers new perspectives and directions for PIV network design.

### v4

@@
 To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR).
-Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics.
-Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness.
-The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
-Abstract
-Particle Image Velocimetry (PIV) is a widely adopted non-invasive imaging technique that tracks the motion of tracer particles across image sequences to capture the velocity distribution of fluid flows.
-It is commonly employed to analyze complex flow structures and validate numerical simulations.
-This study explores the untapped potential of spike cameras—ultra-high-speed, high-dynamic-range vision sensors—in high-speed fluid velocimetry.
-We propose a deep learning framework, Spike Imaging Velocimetry (SIV), tailored for high-resolution fluid motion estimation.
-To enhance the network’s performance, we design three novel modules specifically adapted to the characteristics of fluid dynamics and spike streams: the Detail-Preserving Hierarchical Transform (DPHT), the Graph Encoder (GE), and the Multi-scale Velocity Refinement (MSVR).
-Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), comprising labeled samples from three representative scenarios in fluid dynamics.
-Our proposed method outperforms existing baselines on the PSSD dataset, demonstrating its effectiveness.
-The implementation details of our project, including code and configuration settings, are provided in the supplementary material.
+Furthermore, we introduce a spike-based PIV dataset, Particle Scenes with Spike and Displacement (PSSD), which contains labeled samples from three representative fluid-dynamics scenarios: steady turbulence, high-speed flow, and high-dynamic-range conditions.
+Our proposed method outperforms existing baselines across all these scenarios, demonstrating its effectiveness.
 Introduction
-The need for accurate and non-invasive flow measurement methods has led to the widespread adoption of Particle Image Velocimetry (PIV), a powerful diagnostic tool in fluid dynamics research.
-This technique has been widely applied to critical topics such as studying the aerodynamic shapes of aircraft (Kompenhans et al.
-2000), optimizing vehicle designs to reduce air resistance (Nakagawa et al.
-2016), improving energy conversion efficiency in combustion processes (Chen, Li, and Zhu 2022), understanding the impact of blood flow dynamics on diseases (Raschi et al.
-2012), and predicting the spread of water pollution (MoayeriKashani et al.
-2016).
-Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetS (Cai et al.
-2019b) and RAFT-PIV (Lagemann et al.
-2021).
-In PIV experiments or simulations, tracer particles (e.g., silver-coated particles) are introduced into the fluid to scatter optical signals.
-By estimating their displacements, PIV can infer the velocities of the fluid in which they are suspended.
-The success of optical flow networks (Dosovitskiy et al.
-2015; Sun et al.
-2018; Teed and Deng 2020) has inspired the development of learning-based PIV algorithms (Cai et al.
-2019b, a; Lagemann et al.
-2021).
-Although optical flow networks generally exhibit good generalization ability, the characteristics of fluid scenes cannot be ignored for further performance improvement.
-Turbulence is an irregular flow phenomenon occurring at high Reynolds numbers.
-When measuring the turbulent velocity field with PIV, this scenario is significantly different from macroscopic worlds, with the following key challenges:
-(1) The fluid often moves rapidly within the field of view.
-(2) The scene contains extensive empty regions without tracer particles, yet the network is required to estimate the full flow field.
-(3) Unlike solids, the fluid is highly sparse and unstructured overall, despite the presence of vortices of different scales (called quasi-order structures).
-These challenges often coexist in turbulence.
-Challenge (1) places high demands on the experimental hardware, typically requiring ultra-high-speed cameras.
-Challenge (2) requires addressing potential resolution loss in the network carefully.
-Challenge (3) introduces problems to the extraction of context features.
-To address these challenges, we propose improvements in both hardware and algorithm aspects.
-In terms of hardware, we explore the application of spike cameras (a type of neuromorphic camera) for PIV. We argue that using spike cameras in fluid measurement tasks offers several advantages:
-High temporal resolution (HTR).
-Spike cameras can achieve ultra-high temporal resolution (20000Hz), reducing the interval $\Delta t$ between the source and target time, and providing rich temporal information for accurate predictions.
-This enables spike cameras to handle Challenge (1) effectively;
-High dynamic range (HDR).
-When particles accumulate in large numbers, local regions may become overly bright, exceeding the dynamic range of conventional cameras.
-Spike cameras asynchronously accumulate photons at each pixel, inherently possessing a much larger dynamic range.
-This makes them more advantageous in scenarios with uneven brightness (such as inappropriate lighting or particles moving toward concentrated regions);
-Reduced experimental complexity.
-Spike cameras can continuously capture images at high frame rates with exceptionally low bandwidth and eliminate the need for synchronization controllers to modulate the camera and laser.
+Figure 1: A schematic of fluid velocity field estimation based on spike cameras, and a comparison of the complex velocity field estimated using our proposed SIV network with PIV-NetS (pivnets) and RAFT-PIV (raftpiv).
+The need for accurate and non-invasive flow measurement techniques has led to the widespread use of Particle Image Velocimetry (PIV), a key diagnostic tool for fluid dynamics.
+PIV has been applied to critical areas such as studying aircraft aerodynamics (kompenhans2000particle), optimizing vehicle designs for reduced air resistance (nakagawa2016typical), improving energy efficiency in combustion processes (chen2022investigations), understanding blood flow dynamics in diseases (raschi2012cfd), and predicting water pollution spread (moayerikashani2016tracking).
+In PIV experiments, tracer particles scatter light within a fluid, and their displacements are used to infer the velocity field of the fluid.
+Recent advancements in optical flow networks (flownet; pwcnet; raft) have inspired the development of learning-based PIV algorithms (pivnets; pivliteflownet; raftpiv).
+While optical flow networks are highly generalizable, fluid scenes present unique challenges that limit further performance improvements.
+Turbulence, a chaotic flow phenomenon at high Reynolds numbers, introduces specific issues in velocity field estimation: (1) High-speed fluid movement within the field of view.
+(2) Sparse regions without tracer particles, leading to discrete signals while still requiring the estimation of a continuous flow field.
+(3) Unstructured patterns and shapes with small-scale vortices and quasi-ordered structures.
+These challenges coexist in turbulent flows, necessitating advanced experimental hardware and algorithmic solutions.
+To address these, we propose Spike Imaging Velocimetry (SIV).
+In terms of hardware, we explore the use of spike cameras (firstspike; spikepatent) for PIV.
+Spike cameras offer several advantages for fluid velocity measurement:
+- High Temporal Resolution (HTR): With a temporal resolution of 20,000 Hz, spike cameras significantly reduce the time interval $\Delta t$ between consecutive frames, providing rich temporal information that enables precise predictions.
+This addresses Challenge (1) effectively.
+- High Dynamic Range (HDR): Spike cameras asynchronously accumulate photons at each pixel, providing a larger dynamic range, which is beneficial in scenarios with uneven lighting or high particle concentration.
+This helps mitigate limitations seen in conventional cameras, addressing Challenge (2).
+- Reduced Experimental Complexity: Spike cameras can continuously capture high-speed images with low bandwidth and do not require synchronization controllers, simplifying experimental setups.
 We propose a specialized network, Spike Imaging Velocimetry (SIV), for estimating fluid velocity fields.
@@
 The effectiveness of each proposed module is further validated through ablation studies.
-The main contributions of this work are summarized as follows:
-1.
-We explore a feasible solution for applying spike cameras to fluid velocity measurement, opening new directions for PIV.
+The contributions of this work are summarized as follows:
+1. We explore a feasible solution for applying spike cameras to fluid velocity measurement, opening new directions for PIV.
 2. We design a dedicated network, SIV, tailored to fluid dynamics and turbulent flow properties, and propose three novel modules: the DPHT, GE, and MSVR.
@@
 Spike-based Optical Flow Estimation
-The spike camera (Dong, Huang, and Tian 2021; Huang 2019) is a type of neuromorphic camera (NeurCam).
+The spike camera (firstspike; xiong2024real; su2024intensity) is a type of neuromorphic camera (NeurCam).
 Inspired by biological visual neural structures and mechanisms, NeurCams emulate the encoding principles of the primate retina.
@@
 This process maps the incoming photon stream into the spike stream, which preserves spatiotemporal relationships and can be reconstructed into images.
-For a detailed explanation of the working mechanism, refer to (Dong, Huang, and Tian 2021).
+For a detailed explanation of the working mechanism, refer to (firstspike).
 The application of spike cameras in optical flow estimation provides significant insights for this study.
@@
 Spike-based optical flow estimation is still a relatively new field.
-(Hu et al.
-2022) introduced a specialized neural network architecture for spike-based optical flow estimation.
+(hu2022optical) introduced a specialized neural network architecture for spike-based optical flow estimation.
 This structure can adaptively eliminate motion blur in spike streams based on prior motion information.
-A framework based on RAFT is proposed by (Zhao et al.
-2022) to handle continuous spike streams.
-(Zhao et al.
-2024) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network.
-(Chen, Yu, and Huang 2023) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by (Xia et al.
-2024).
+A framework based on RAFT is proposed by (spike2flow) to handle continuous spike streams.
+(hist) demonstrates that using a multi-layer hierarchical structure to characterize spike streams can effectively suppress the influence of randomness in spikes and enhance the feature extraction capability of the network.
+(chen2023self) proposed a self-supervised approach to learn optical flow from spike cameras, which is then extended by (xia2024unsupervised).
 Image-based PIV Algorithms
-Many traditional image-based PIV algorithms (Lourenco and Krothapalli 1995; Nobach and Honkanen 2005) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation windows.
+Many traditional image-based PIV algorithms (classicpiv1; classicpiv3) usually involve applying a fast Fourier transform (FFT) between two subregions of image pairs, known as interrogation windows.
 Displacement is then determined by finding the maximum correlation between the two interrogation windows, which is similar to the correlation method in optical flow estimation to some extent.
 However, traditional methods depend on finely tuned parameters and complex optimization techniques.
-With advancements in optical flow networks, various architectures (Dosovitskiy et al.
-2015; Sun et al.
-2018; Teed and Deng 2020) have been developed to predict dense velocity fields of macroscopic objects with high accuracy.
-Early PIV networks like PIV-DCNN (Lee, Yang, and Yin 2017) laid the groundwork for these developments.
-PIV-NetS (Cai et al.
-2019b) is capable of generating full-resolution displacement outputs.
-(Cai et al.
-2019a) adapt LiteFlowNet (Hui, Tang, and Loy 2018) to PIV.
-RAFT-PIV (Lagemann et al.
-2021), derived from RAFT (Teed and Deng 2020), has shown exceptional performance on multiple datasets.
-(Yu et al.
-2021) removes the computationally inefficient components of RAFT, such as the context encoder, to improve inference speed.
-Additionally, unsupervised learning frameworks for PIV(Zhang and Piggott 2020; Zhang et al.
-2023a) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements.
-Other recent PIV networks include LIMA (Manickathan, Mucignat, and Lunati 2023), CC-FCN (Gao et al.
-2021), En-FlowNetC (Ji, An, and Hui 2024), ARAFT-FlowNet (Han and Wang 2023), OFVNetS (Ji, Hui, and An 2024), and PIV-PWCNet (Zhang et al.
-2023b).
+With advancements in optical flow networks, various architectures (flownet; pwcnet; raft) have been developed to predict dense velocity fields of macroscopic objects with high accuracy.
+Early PIV networks like PIV-DCNN (pivdcnn) laid the groundwork for these developments.
+PIV-NetS (pivnets) is capable of generating full-resolution displacement outputs.
+(pivliteflownet) adapt LiteFlowNet (liteflownet) to PIV.
+RAFT-PIV (raftpiv), derived from RAFT (raft), has shown exceptional performance on multiple datasets.
+(lightpivnet) removes the computationally inefficient components of RAFT, such as the context encoder, to improve inference speed.
+Additionally, unsupervised learning frameworks for PIV(unsupervisedpiv1; unsupervisedpiv2) have been introduced, reducing reliance on the extremely difficult-to-obtain ground-truth displacements.
+Other recent PIV networks include LIMA (manickathan2023lightweight), CC-FCN (gao2021robust), En-FlowNetC (ji2024cross), ARAFT-FlowNet (han2023attention), OFVNetS (ji2024effects), and PIV-PWCNet (zhang2023pyramidal).
 However, the performance gains of most PIV networks are largely attributed to advances in optical flow estimation architectures, rather than targeted enhancements tailored to the specific characteristics of fluid dynamics, and most implementations remain unavailable as open-source.
 Event-based PIV Algorithms
-The event camera (Lichtsteiner, Posch, and Delbruck 2008; Moeys et al.
-2017; Posch, Matolin, and Wohlgenannt 2010; Huang, Guo, and Chen 2017) is another type of NeurCam, which is inspired by the peripheral retina (Rebecq et al.
-2019).
+The event camera (dvs; davis; atis; jiang2024edformer) is another type of NeurCam, which is inspired by the peripheral retina (event1).
 Unlike spike cameras, event cameras independently monitor the change in light intensity (increase or decrease) at each pixel and generate signals based on the local pixel information of dynamic changes.
-When the brightness change of a pixel exceeds a preset threshold, the camera records this change event (brightness increase as a positive event, brightness decrease as a negative event) and outputs its timestamp, position, and change value as data.
-However, due to the lack of absolute intensity values, event cameras struggle to reconstruct complete images, particularly in static regions and with fine textures.
-The application of event cameras to PIV has already been explored by (Willert and Klinner 2022) and (Willert 2023).
-Similar to the spike camera, the event camera offers higher temporal resolution, high dynamic range, low latency, and reduced experimental complexity.
+When the intensity change of a pixel exceeds a preset threshold, the camera records this change event (intensity increase as a positive event, intensity decrease as a negative event) and outputs its timestamp, position, and change value as data.
+The success of event cameras in optical flow estimation (lee2020spike-flownet; shiba2022secrets) has paved the way for their application to PIV.
+(EBIV) introduced the first event-camera PIV algorithm based on correlation and motion compensation.
+(eventpiv) combined an event camera with pulsed illumination, substantially reducing estimation latency.
+These studies are insightful and practical; however, their algorithmic novelty is limited, and robustness is partly constrained by the event camera’s lack of absolute intensity.
 Preliminary
@@
 The Integral-and-Fire (IF) mechanism of spike cameras can be mathematically represented as follows:
-$A(\mathbf{x},t)=\int_{0}^{t}\alpha\cdot I(\mathbf{x},\tau)\bmod\theta,$ (1)
+$A(\mathbf{x},t)=\left(\int_{0}^{t}\alpha\cdot I(\mathbf{x},\tau)d\tau\right)\bmod\theta,$ (1)
 Where $A(\mathbf{x},t)$ represents the voltage of the accumulator at pixel $\mathbf{x}=(x,y)$ at time $\tau$, and $\theta$ is the threshold of the comparator.
 The spike camera’s output is a spatiotemporal binary stream $S$ of size $H\times W\times N$, where $H$ and $W$ are the height and width of the sensor, respectively, and $N$ represents the temporal dimension.
-Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and brightness representation.
+Compared to traditional cameras, spike cameras preserve richer visual information regarding both temporal resolution and intensity representation.
 Data Generation
+Training neural networks using supervised strategies requires data with ground truth to optimize model parameters.
+However, it is difficult to obtain accurate velocity fields from experimental PIV data.
+We have to generate a synthetic dataset for training, following (pivnets).
 Our proposed dataset, Particle Scenes with Spike and Displacement (PSSD), includes three sub-datasets: Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3).
@@
 We use the reference flow field to simulate the motion of the scene and construct a virtual spike camera to trigger the spike stream.
-The reference flow field data are taken from the high-precision numerical simulation database used for turbulence research, Johns Hopkins Turbulence Database (JHTDB) (Li et al.
-2008).
-JHTDB employs Direct Numerical Simulation (DNS) to simulate turbulent flow fields.
-DNS directly solves the Navier-Stokes equations without any model assumptions or simplifications, enabling precise capture of all scales in turbulent flow.
-This method resolves the fluid motion in both space and time with high-resolution grids, accurately capturing the complex and multi-scale nature of turbulence.
-Additional details about the dataset can be found in our supplementary materials.
+Both the images and the spike stream in the dataset include added noise.
+The reference flow fields are obtained from the Johns Hopkins Turbulence Database (JHTDB) (JHTDB), a high-precision numerical simulation database widely used in turbulence research.
+JHTDB is based on direct numerical simulation (DNS) of the Navier–Stokes equations without any turbulence modeling, resolving all dynamically relevant scales in both space and time on fine grids.
+As a result, the simulated flow fields closely approximate real turbulent flows and can be regarded as ground truth in our study.
+Figure 2: (a) Normal illumination scene (Problem 1, Problem 2); (b) High dynamic range scene (Problem 3).
 Approaches
@@
 At the pixel level, for any given pixel $\mathbf{p}=(x,y)$ at $t$, we need to find a unique correspondence formulated by:
-$\mathcal{I}(\mathbf{p},t)\leftarrow\mathcal{I}(\mathbf{p},t+\Delta t).$ (2)
+$\mathcal{I}(\mathbf{p},t)\leftarrow\mathcal{I}(\mathbf{p}+\mathbf{D}(\mathbf{p}),t+\Delta t).$ (2)
 Here, $\mathcal{I}$ is the mapping of the spike stream $\mathbf{S}$ in the high-dimensional feature space, representing the scene information.
@@
 Detail-Preserving Hierarchical Transform
-In order to avoid the loss of particle signal information during feature extraction, we propose a multi-scale pyramidal architecture, named Detail-Preserving Hierarchical Transform (DPHT), which incorporates a 3D Detail-Preserving Downsampling module (as shown in Figure 3(a)).
-3D Detail-Preserving downsampling. As shown in Figure 3(a), suppose the input of level $l$ is a feature vector $\mathbf{x}^{(l)}$, and the processing of $\mathbf{x}^{(l)}$ at this level is as follows:
+To preserve particle signal information during feature extraction, we propose the Detail-Preserving Hierarchical Transform (DPHT), a multi-scale pyramidal architecture that integrates a 3D Detail-Preserving Downsampling module (see Figure 4(a)).
+3D Detail-Preserving Downsampling. Given an input feature vector $\mathbf{x}^{(l)}$ at level $l$, the processing is performed as follows:
 $\mathbf{x}^{(l)}_{\mathrm{out}}=\mathscr{D}^{(l)}(\mathbf{x}^{(l)})=\mathscr{F}_{\mathrm{2C}}(\mathscr{F}_{\mathrm{3C}}(\mathbf{x}^{(l)})+\mathscr{F}_{\mathrm{3MP}}(\mathbf{x}^{(l)})),$ (3)
-Here, $\mathscr{F}_{\mathrm{3C}}$ represents a 5×3×3 3D convolution, $\mathscr{F}_{\mathrm{3MP}}$ is a combination of a 3D max pooling operation and a 5×1×1 3D convolution, and $\mathscr{F}_{\mathrm{2C}}$ refers to a sequential stack of 2D convolutions.
-The output of the previous level serves as the input to the next level, i.e., $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}_{\mathrm{out}}=\mathscr{D}^{(l)}(\mathbf{x}^{(l)})$.
-Global temporal aggregation.
-In this step, the output feature vectors from each level are further refined by several convolutions (for detailed information, refer to the supplementary materials), and then concatenated together to form the final representation of the DPHT:
-$\mathbf{R}=\mathscr{A}_{m}[\mathbf{Concat}(\mathbf{x}^{(l=1)}_{\mathrm{out}},\mathbf{x}^{(l=2)}_{\mathrm{out}},\mathbf{x}^{(l=3)}_{\mathrm{out}})$], (4)
-$\mathscr{A}_{m}$ denotes convolution. After processing with DPHT, the signals of the source and target sub-streams are denoted as $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$, respectively.
+where $\mathscr{F}_{\mathrm{3C}}$ is a 5×3×3 3D convolution, $\mathscr{F}_{\mathrm{3MP}}$ combines 3D max pooling and a 5×1×1 3D convolution, and $\mathscr{F}_{\mathrm{2C}}$ refers to a stack of 2D convolutions.
+The output of each level is passed to the next, i.e., $\mathbf{x}^{(l+1)}=\mathbf{x}^{(l)}_{\mathrm{out}}$.
+Global Temporal Aggregation. The output feature vectors from each level are refined through additional convolutions and then concatenated to form the final DPHT representation:
+$\mathbf{R}=\mathscr{A}{m}[\mathbf{Concat}(\mathbf{x}^{(l=1)}_{\mathrm{out}},\mathbf{x}^{(l=2)}_{\mathrm{out}},\mathbf{x}^{(l=3)}_{\mathrm{out}})$], (4)
+where $\mathscr{A}_{m}$ denotes a convolution operation. The source and target sub-streams’ signals are represented as $\mathbf{R}_{s}$ and $\mathbf{R}_{t}$, respectively.
 Graph Encoder
-When the Reynolds number is high, fluid flow becomes turbulent.
-As energy cascades and dissipates, large vortices become unstable and gradually break into smaller ones until they vanish.
-This indicates that fluid motion is not purely chaotic, but rather exhibits an underlying topological structure.
-To capture the contextual information of the fluid, we propose a Graph Encoder (GE).
-It is a feature extraction network based on Graph Neural Networks (GNNs) and Convolutional Neural Networks (CNNs).
-By representing features as graph nodes and constructing a dynamic adjacency matrix, GE flexibly aggregates and updates information from neighboring nodes, enabling a larger receptive field than conventional CNNs.
-Graph Attention Networks (GATs) (Veličković et al.
-2017), leveraging attention mechanisms, learn the relative contribution of each neighbor to a node’s new representation, thereby mitigating interference from particle-sparse regions.
-The serial combination of GNN and CNN enables adaptive global perception and local refinement, effectively capturing fluid-specific contextual features.
+In turbulent flows with high Reynolds numbers, large vortices break into smaller ones, revealing an underlying topological structure in fluid motion.
+To capture this structure, we propose the Graph Encoder (GE), a feature extraction network based on Graph Neural Networks (GNNs) and Convolutional Neural Networks (CNNs).
+By representing features as graph nodes and constructing a dynamic adjacency matrix, GE aggregates and updates information from neighboring nodes, allowing for a larger receptive field than traditional CNNs.
+Graph Attention Networks (GATs) (GAT) leverage attention mechanisms to learn the relative contribution of each neighbor, reducing interference from particle-sparse regions.
+This combination of GNN and CNN enables adaptive global perception and local refinement, effectively capturing fluid-specific contextual features.
 Graph Projection.
-The first step is to map $\mathbf{R}_{s}$ extracted from DPHT from the general coordinate space to the node space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{s})$.
-The mapped context node representations, denoted as $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}\in\mathbb{R}^{C\times K}$, contain appearance features related to the shape and regional information of the scene.
-$C$ indicates channel number and $K$ is a hyper-parameter representing the number of nodes.
-In this paper, we set $K=128$.
-$\mathscr{P}_{R\to V}(\cdot)$ is a linear combination of feature vectors within the grid space, formulated by,
-$v_{c,i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{s,j}).$ (5)
-Here, $\mathscr{N}(\cdot)$ is the $L_{2}$-normalization function applied to the channel dimension of each node vector, and $z_{ij}$ denotes the projection weights computed by a CNN, i.e., $z=\mathscr{F}_{R\to V}(\mathbf{R}_{s})\in\mathbb{R}^{(H\times W)\times K}$.
+The first step is to map $\mathbf{R}_{s}$ from the DPHT to the node space, i.e., $\mathbf{V}_{c}=\mathscr{P}_{R\to V}(\mathbf{R}_{s})$.
+The context node representations, $\mathbf{V}_{c}=\{v_{c}^{1},\cdots,v_{c}^{n}\}\in\mathbb{R}^{C\times K}$, contain appearance features related to the scene’s shape and regional information, where $C=128$ is the channel number and $K$ is the number of nodes (set to $K=128$ in this paper).
+The projection is formulated as:
+$v_{c}^{i}=\mathscr{N}(\sum\limits_{\forall{j}}z_{ij}\cdot\mathbf{R}_{s,j}),$ (5)
+where $\mathscr{N}(\cdot)$ is the $L_{2}$ normalization, and $z_{ij}$ denotes the normalized projection weights computed by a $1\times 1$ convolution and a softmax function.
 Graph Convolution Layer.
-After obtaining the nodes $\mathbf{v}_{c}$ in the graph space, the dynamic adjacency matrix $\widetilde{\mathbf{A}}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ can be derived by calculating the similarity between all node vectors.
-The detailed definition of GAT can be found in the supplementary material.
-Here, we implicitly denote the operations of the two GAT modules as $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$.
-The updated node features are given by
+After obtaining the nodes $\mathbf{v}_{c}$, the dynamic adjacency matrix $\widetilde{\mathbf{A}}=\mathbf{V}_{c}^{T}\mathbf{V}_{c}$ is computed based on the similarity between node vectors.
+The operations of two GAT modules, $\mathscr{G}_{1}$ and $\mathscr{G}_{2}$, update the node features as follows:
 $\mathbf{V_{c}^{\mathrm{out}}}=\mathscr{G}_{2}[\mathscr{G}_{1}(\mathbf{V}_{c},\widetilde{\mathbf{A}})+\mathbf{V}_{c},\widetilde{\mathbf{A}}]+\mathscr{G}_{1}(\mathbf{V}_{c},\widetilde{\mathbf{A}})+\mathbf{V}_{c},$ (6)
 Graph Re-projection.
-Our graph re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}=[z_{ij}]$ as inputs, generating a 2D feature map $\widetilde{\mathbf{R}}_{c}$.
-Since inverting the projection weight matrix is not feasible, we follow the approach in (Li and Gupta 2018; Luo et al.
-2022) and compute the enhanced grid context features $\widetilde{\mathbf{R}}_{c}$ by re-weighting the vertex features $\mathbf{V_{c}^{\mathrm{out}}}$.
-This is given by,
-$\widetilde{\mathbf{R}}_{c}=\mathscr{P}_{V\to R}(\mathbf{V_{c}^{\mathrm{out}}},\mathbf{Z})=\mathbf{Z}\mathbf{V_{c}^{\mathrm{out}}}^{T}.$ (7)
-Thus, $\mathscr{P}_{V\to R}$ linearly interpolates 2D pixel features based on the node assignments of the 2D grid features, without additional parameters.
-Note that even if two pixels are assigned to the same node, they will have different features after re-projection.
-Therefore, the GCU may preserve the spatial details of the signal.
-Finally, these projection results can be integrated into a residual extractor network $\mathscr{R}$ for local refinement.
-Final context feature $\mathbf{F}_{c}$ given by,
+The re-projection operation $\mathscr{P}_{V\to R}(\cdot)$ takes the transformed vertex features $\mathbf{V_{c}^{\mathrm{out}}}$ and the projection weight matrix $\mathbf{Z}$, producing a 2D feature map $\widetilde{\mathbf{R}}c$:
+$\widetilde{\mathbf{R}}_{c}=\mathscr{P}_{V\to R}(\mathbf{V_{c}^{\mathrm{out}}},\mathbf{Z})=\mathbf{Z}\left(\mathbf{V_{c}^{\mathrm{out}}}\right)^{T}.$ (7)
+This operation preserves spatial details, as even if two pixels share the same node, they will have distinct features.
+Finally, these results are integrated into a residual extractor network $\mathscr{R}$ for local refinement:
 $\mathbf{F}_{c}=\mathscr{R}(\mathbf{R}_{s}+\alpha\widetilde{\mathbf{R}}_{c}),$ (8)
-where $\alpha$ denotes a learnable parameter that is initialized as 0 and gradually performs a weighted sum.
+where $\alpha$ is a learnable parameter initialized to 0, gradually refining the feature sum.
 Multi-scale Iterative Optimizer
-We adopt the Multi-scale Iterative Optimizer (MSIO) from DIP (Zheng et al.
-2022) as the backbone for iterative residual flow estimation in SIV.
+We adopt the Multi-scale Iterative Optimizer (MSIO) from DIP (dip) as the backbone for iterative residual flow estimation in SIV.
 Unlike RAFT’s full correlation volume construction, MSIO leverages Inverse PatchMatch to significantly reduce the computational cost introduced by large-scale feature maps (1/4 resolution of the full size).
-The detailed architecture of MSIO can be found in (Zheng et al.
-2022); here, we denote it implicitly as $\mathscr{U}$.
+The detailed architecture of MSIO can be found in (dip); here, we denote it implicitly as $\mathscr{U}$.
 The output multi-scale velocity field is given by
-$\{\mathbf{Vel}^{i}\}_{i=1}^{\text{iters}}=\mathscr{U}(\mathbf{F}_{c},\mathbf{F}_{m}),$ (9)
-$\mathbf{F}_{m}$ denotes the motion features extracted by the Motion Encoder.
-iter denotes the number of iterations of GRU (Teed and Deng 2020).
-The detailed architecture of the Motion Encoder is provided in the supplementary material.
+$\{\mathbf{u}_{i}\}_{i=1}^{N}=\mathscr{U}(\mathbf{F}_{c},\mathbf{F}_{m}),$ (9)
+where $\mathbf{F}_{m}$ denotes motion features extracted by the Motion Encoder, and $N$ refers to the number of GRU iterations (raft).
 Multi-scale Velocity Refinement
+Figure 5: Multi-scale Velocity Refinement Module.
 DIP adopts the same compact and effective convex upsampling strategy as RAFT to recover full-resolution optical flow.
@@
 To address this, we propose a Multi-scale Velocity Refinement (MSVR) module to assist in reconstructing the complete velocity field, including fine-scale vortices.
-As illustrated in Figure 4, we first apply a sequence of 2D convolutions followed by activation functions to process the multi-scale velocity fields generated by MCIO.
+As illustrated in Figure 5, we first apply a sequence of 2D convolutions followed by activation functions to process the multi-scale velocity fields generated by MSIO.
 Information exchange across scales is achieved via cross convolution from coarse to fine levels.
-The extracted features are then fused and processed by distinct convolutional networks to generate a residual velocity field $\mathbf{Vel}^{\text{res}}$ and a quality map $\mathbf{Q}$.
+The extracted features are then fused and processed by distinct convolutional networks to generate a residual velocity field $\mathbf{u}_{\text{res}}$ and a quality map $\mathbf{Q}$.
 The refined fluid velocity field is given by,
-$\mathbf{Vel}^{*}=\mathbf{Vel}^{i=\text{iters}}+\mathbf{Vel}^{\text{res}}\odot\mathbf{Q},$ (10)
-$\odot$ denotes element-wise multiplication.
-Figure 4: Multi-scale Velocity Refinement Module.
+$\mathbf{u}_{\text{ref}}=\mathbf{u}_{N}+\mathbf{u}_{\text{res}}\odot\mathbf{Q},$ (10)
+where $\odot$ denotes element-wise multiplication.
 Loss Function
-Since SIV outputs a prediction sequence of length $N$ at each iteration, the flow loss is defined as:
-$L_{\text{flow}}=\sum_{i=0}^{i=N}w_{i}\cdot\|f_{i}-f_{gt}\|_{1}.$ (11)
-We introduce a gradient loss to prevent the generated flow field from becoming overly smooth in regions like shear layers or small vortical structures. It is computed as follows:
-$L_{\text{grad}}=\sum_{i=0}^{i=N}w_{i}\cdot(\|\nabla_{\mathbf{x}}f_{i}-\nabla_{\mathbf{x}}f_{gt}\|_{1}+\|\nabla_{\mathbf{y}}f_{i}-\nabla_{\mathbf{y}}f_{gt}\|_{1}).$ (12)
-We set $w_{i}=\gamma^{N-i-1}$ and $\gamma=0.8$ for both $L_{\text{flow}}$ and $L_{\text{grad}}$. The total loss can be formulated as follows:
+The loss function for the proposed network is composed of two parts: the flow loss and the gradient loss.
+Suppose the recurrent optimizer of the network has $N$ iterations, and the estimated flow fields of each iteration are $\{\mathbf{u}_{1},\dots,\mathbf{u}_{N},\mathbf{u}_{N+1}\}$, $\mathbf{u}_{N+1}=\mathbf{u}_{\text{ref}}$.
+The flow loss is defined as:
+$L_{\text{flow}}=\sum_{i=1}^{N+1}\gamma^{N+1-i}\cdot\|\mathbf{u}_{i}-\mathbf{u}_{\text{gt}}\|_{1},$ (11)
+$\mathbf{u}_{\text{gt}}$ is the ground truth of velocity field. To preserve small vortical structures, we introduce a gradient loss:
+$L_{\text{grad}}$ $=\sum_{i=1}^{N+1}\gamma^{N+1-i}\cdot(\|\nabla_{\mathbf{x}}\mathbf{u}_{i}-\nabla_{\mathbf{x}}\mathbf{u}_{\text{gt}}\|_{1}$ (12)
+$+\|\nabla_{\mathbf{y}}\mathbf{u}_{i}-\nabla_{\mathbf{y}}\mathbf{u}_{\text{gt}}\|_{1}).$
+The total loss is:
 $L=L_{\text{flow}}+\beta\cdot L_{\text{grad}}.$ (13)
-Here, we set $\beta=0.3$.
-Experimental Results
-In this section, we evaluate our method on the PSSD dataset and compare its performance with other traditional and learning-based methods.
-Additionally, we conduct a series of ablation experiments to assess the effectiveness of the proposed modules.
+$\gamma$ is set to 0.8, and $\beta$ is set to 0.3.
+Experiments
+We evaluate our method on the PSSD dataset and compare its performance with both traditional and learning-based approaches.
+Additionally, we conduct a series of ablation studies to assess the effectiveness of the proposed modules.
 Figure 6: Visualization examples of different algorithms in Steady Turbulence (Problem 1), High-speed Flow (Problem 2), and HDR Scenes (Problem 3) are shown.
@@
 SIV is implemented using the PyTorch framework on a single NVIDIA RTX 2080 Ti GPU.
-We employ the Adam optimizer with a warm-up strategy during training.
-For each sub-dataset, we train the model for 100 epochs with an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs.
-The number of iterative refinements of GRUs is set to 12.
-The batch size is set to 4.
-Following (Cai et al.
-2019b) and (Lagemann et al.
-2021), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms.
-Additionally, we incorporate another widely used PIV metric, Root Mean Squared Error (RMSE), in our extensive experiments, as detailed in the supplementary material.
+We use the Adam optimizer with a warm-up strategy during training.
+For each sub-dataset, the model is trained for 100 epochs with a batch size of 4 and an initial learning rate of $10^{-4}$, applying an 85% decay every 10 epochs.
+The number of iterative refinements for the GRUs is set to 12.
+Following (pivnets) and (raftpiv), we use endpoint error (EPE) as the primary evaluation metric for comparing different algorithms.
+The EPE is the spatially mean value of Euclidean distance between the predicted flow $\mathbf{u}_{\text{pred}}$ and its ground truth $\mathbf{u}_{\text{gt}}$.
 Methods Channel Isotropic Mhd Mixing Average
 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11
-coarse-to-fine HS 6.790 0.509 7.415 1.998 7.683 2.558 9.388 1.954 7.819 1.755
-RAFT-PIV-Image 0.377 0.215 2.181 1.252 1.336 0.842 1.875 1.064 1.442 0.843
-RAFT-PIV-Spike 0.347 0.208 1.223 0.755 0.857 0.557 0.958 0.581 0.846 0.525
-PIV-NetS-Image 0.545 0.301 2.560 1.497 1.574 0.985 2.247 1.320 1.731 1.026
-PIV-NetS-Spike 0.553 0.289 1.507 0.874 1.252 0.734 1.382 0.782 1.173 0.670
-SCFlow 0.354 0.222 2.232 1.305 1.374 0.901 2.132 1.169 1.522 0.899
-HiST-SFlow 0.369 0.224 1.271 0.796 0.964 0.611 1.029 0.637 0.908 0.567
-Flowformer-Spike 0.286 0.147 1.069 0.642 0.727 0.437 0.806 0.472 0.722 0.425
-SIV (ours) 0.239 0.159 0.896 0.587 0.613 0.422 0.680 0.441 0.607 0.402
-Table 1: EPE ($\downarrow$) results on PSSD dataset (Problem 1). The best results are marked in bold.
-Comparative Experiments
-We compare our method against several open-source baselines, including a classical PIV algorithm (Cai et al.
-2018), PIV networks (PIV-NetS (Cai et al.
-2019b), RAFT-PIV (Lagemann et al.
-2021)), spike-adapted general optical flow networks (e.g., Flowformer with spike input (Huang et al.
-2022)), and spike-based optical flow networks (SCFlow (Hu et al.
-2022), HiST-SFlow (Zhao et al.
-2024)).
+(A) coarse-to-fine HS 6.790 0.509 7.415 1.998 7.683 2.558 9.388 1.954 7.819 1.755
+(B) RAFT-PIV-Image 0.377 0.215 2.181 1.252 1.336 0.842 1.875 1.064 1.442 0.843
+(C) RAFT-PIV-Spike 0.347 0.208 1.223 0.755 0.857 0.557 0.958 0.581 0.846 0.525
+(D) PIV-NetS-Image 0.545 0.301 2.560 1.497 1.574 0.985 2.247 1.320 1.731 1.026
+(E) PIV-NetS-Spike 0.553 0.289 1.507 0.874 1.252 0.734 1.382 0.782 1.173 0.670
+(F) SCFlow 0.354 0.222 2.232 1.305 1.374 0.901 2.132 1.169 1.522 0.899
+(G) HiST-SFlow 0.369 0.224 1.271 0.796 0.964 0.611 1.029 0.637 0.908 0.567
+(H) Flowformer-Spike 0.286 0.147 1.069 0.642 0.727 0.437 0.806 0.472 0.722 0.425
+(I) SIV (ours) 0.239 0.159 0.896 0.587 0.613 0.422 0.680 0.441 0.607 0.402
+Table 1: EPE ($\downarrow$) results on PSSD dataset - Problem 1. The best results are marked in bold.
 Methods Channel Isotropic Mhd Mixing Average
 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11
-coarse-to-fine HS 20.691 8.244 17.164 8.051 17.469 8.170 17.131 7.936 18.114 8.100
-RAFT-PIV-Image 0.665 0.368 4.730 2.635 2.320 1.366 4.305 2.368 3.005 1.684
-RAFT-PIV-Spike 0.688 0.412 2.410 1.483 1.468 0.959 2.071 1.241 1.659 1.024
-PIV-NetS-Image 1.163 0.683 4.866 2.830 2.673 1.643 4.347 2.500 3.262 1.914
-PIV-NetS-Spike 0.887 0.540 3.172 1.818 2.392 1.410 3.067 1.744 2.379 1.378
-SCFlow 0.694 0.403 4.689 2.613 2.387 1.391 4.326 2.380 3.024 1.696
-HiST-SFlow 0.556 0.364 2.212 1.365 1.404 0.998 1.911 1.179 1.521 0.976
-Flowformer-Spike 0.533 0.284 1.975 1.179 1.214 0.748 1.668 0.972 1.348 0.796
-SIV (ours) 0.370 0.254 1.687 1.080 1.071 0.722 1.449 0.912 1.144 0.742
-Table 2: EPE ($\downarrow$) results on PSSD dataset (Problem 2). The best results are marked in bold.
-As shown in Table 1, 2, and 3, $\Delta t=21$ and $\Delta t=11$ represent the number of timestamps for the central temporal intervals between two spike streams used for velocity field prediction, respectively.
-For detailed descriptions, please refer to the supplementary material.
-Our proposed SIV method achieves the lowest average End-Point Error (EPE) across all evaluated sub-datasets.
-Compared to baseline approaches, SIV demonstrates superior prediction accuracy in both low-velocity (Problem 1) and high-velocity (Problem 2) optical flow scenarios.
-Furthermore, in high dynamic range conditions (Problem 3), SIV effectively exploits the inherent advantages of spike-based input modalities to address the fundamental limitations of conventional frame-based imaging systems.
-Figure 5 presents the visualized predictions of all methods across three datasets.
-The top row shows the velocity fields visualized using the standard optical flow color coding (Baker et al.
-2011).
-The bottom row shows the endpoint error (EPE) maps between the predicted and ground-truth flow.
-Shear layers are often a primary source of turbulence generation and development.
-SIV achieves the lowest prediction error in both shear layers and small-scale vortices, while avoiding noticeable blurring, distortion, or discontinuities.
+(A) coarse-to-fine HS 20.691 8.244 17.164 8.051 17.469 8.170 17.131 7.936 18.114 8.100
+(B) RAFT-PIV-Image 0.665 0.368 4.730 2.635 2.320 1.366 4.305 2.368 3.005 1.684
+(C) RAFT-PIV-Spike 0.688 0.412 2.410 1.483 1.468 0.959 2.071 1.241 1.659 1.024
+(D) PIV-NetS-Image 1.163 0.683 4.866 2.830 2.673 1.643 4.347 2.500 3.262 1.914
+(E) PIV-NetS-Spike 0.887 0.540 3.172 1.818 2.392 1.410 3.067 1.744 2.379 1.378
+(F) SCFlow 0.694 0.403 4.689 2.613 2.387 1.391 4.326 2.380 3.024 1.696
+(G) HiST-SFlow 0.556 0.364 2.212 1.365 1.404 0.998 1.911 1.179 1.521 0.976
+(H) Flowformer-Spike 0.533 0.284 1.975 1.179 1.214 0.748 1.668 0.972 1.348 0.796
+(I) SIV (ours) 0.370 0.254 1.687 1.080 1.071 0.722 1.449 0.912 1.144 0.742
+Table 2: EPE ($\downarrow$) results on PSSD dataset - Problem 2. The best results are marked in bold.
 Methods Channel Isotropic Mhd Mixing Average
 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11 $\Delta t$=21 $\Delta t$=11
-coarse-to-fine HS 7.908 1.044 7.924 2.706 8.345 3.130 7.854 2.716 8.008 2.399
-RAFT-PIV-Image 0.447 0.244 2.694 1.495 1.447 0.888 2.311 1.271 1.725 0.974
-RAFT-PIV-Spike 0.320 0.200 1.203 0.740 0.798 0.530 0.946 0.576 0.817 0.512
-PIV-NetS-Image 0.667 0.343 2.786 1.602 1.641 0.956 2.393 1.378 1.872 1.070
-PIV-NetS-Spike 0.476 0.247 1.420 0.856 1.157 0.716 1.342 0.774 1.099 0.648
-SCFlow 0.462 0.269 1.850 0.837 1.300 0.774 1.972 0.769 1.396 0.662
-HiST-SFlow 0.329 0.180 1.179 0.640 0.832 0.452 0.960 0.476 0.825 0.437
-Flowformer-Spike 0.269 0.145 1.052 0.620 0.716 0.429 0.813 0.481 0.713 0.419
-SIV (ours) 0.234 0.155 0.880 0.573 0.598 0.404 0.677 0.438 0.598 0.392
-Table 3: EPE ($\downarrow$) results on PSSD dataset (Problem 3). The best results are marked in bold.
+(A) coarse-to-fine HS 7.908 1.044 7.924 2.706 8.345 3.130 7.854 2.716 8.008 2.399
+(B) RAFT-PIV-Image 0.447 0.244 2.694 1.495 1.447 0.888 2.311 1.271 1.725 0.974
+(C) RAFT-PIV-Spike 0.320 0.200 1.203 0.740 0.798 0.530 0.946 0.576 0.817 0.512
+(D) PIV-NetS-Image 0.667 0.343 2.786 1.602 1.641 0.956 2.393 1.378 1.872 1.070
+(E) PIV-NetS-Spike 0.476 0.247 1.420 0.856 1.157 0.716 1.342 0.774 1.099 0.648
+(F) SCFlow 0.462 0.269 1.850 0.837 1.300 0.774 1.972 0.769 1.396 0.662
+(G) HiST-SFlow 0.329 0.180 1.179 0.640 0.832 0.452 0.960 0.476 0.825 0.437
+(H) Flowformer-Spike 0.269 0.145 1.052 0.620 0.716 0.429 0.813 0.481 0.713 0.419
+(I) SIV (ours) 0.234 0.155 0.880 0.573 0.598 0.404 0.677 0.438 0.598 0.392
+Table 3: EPE ($\downarrow$) results on PSSD dataset - Problem 3. The best results are marked in bold.
+Comparative Experiments
+We compare our method with several open-source baselines, including classical PIV algorithms (cai2018motion), image-based PIV networks (PIV-NetS (pivnets), RAFT-PIV (raftpiv)), spike-adapted optical flow networks (e.g., Flowformer with spike input (flowformer)), and spike-based optical flow networks (SCFlow (hu2022optical), HiST-SFlow (hist)).
+All networks trained on spikes (incl.
+(C)(E)(F)(G)(H)(I) in Table 1, 2, and 3) use voxel-grid inputs to ensure a fair evaluation.
+All image baselines (incl.
+(A)(B)(D) in Table 1, 2, and 3) use sharp frames containing only artificially added noise (as shown in Figure 2, not spike reconstructions) simulating high-speed cameras to highlight the differences between image and spike inputs.
+The image is temporally aligned with the intermediate moments of the spike stream.
+As shown in Tables 1, 2, and 3, $\Delta t=21$ and $\Delta t=11$ represent the time intervals between two spike streams used for velocity field prediction.
+Our SIV method achieves the lowest average EPE across all sub-datasets.
+Compared to baseline approaches, SIV demonstrates superior prediction accuracy in both low-velocity (Problem 1) and high-velocity (Problem 2) scenarios.
+Additionally, in high dynamic range conditions (Problem 3), SIV effectively leverages the advantages of spike-based input modalities, addressing the fundamental limitations of conventional frame-based systems.
+Figure 6 visualizes the predictions on three problems.
+The top row visualizes the velocity fields using the standard optical flow color coding (colorcoding), and the bottom row shows the EPE.
+Since shear layers are crucial for turbulence generation and development, it is noteworthy that SIV attains the lowest prediction error in both shear layers and small-scale vortices, while avoiding noticeable blurring, distortion, or discontinuities.
 Ablation Study
 Ablation of modules.
-We conduct a series of ablation experiments to evaluate the contribution of each component to overall performance, as detailed in Table 4.
-All the values are the arithmetic mean of the scenes.
-The best results are marked in bold.
-These results underscore the effectiveness of the DPHT, GE, MSVR and $L_{\text{grad}}$.
-In Table 4, the comparison between experiments (A) and (B) demonstrates the effectiveness of DPHT.
-Both the comparison between (B) and (C) and that between (D) and (E) demonstrate the effectiveness of the Graph Encoder (GE).
-Both the comparison between (B) and (D) and that between (C) and (E) highlight the contribution of MSVR.
+All values in Table 4 and 5 are the arithmetic mean across scenes ($\Delta t=21$), with the best results highlighted in bold.
+The results emphasize the effectiveness of DPHT, GE, MSVR, and $L_{\text{grad}}$.
+In Table 4, comparisons between experiments (A) and (B) show the effectiveness of DPHT, while comparisons between (B) and (C), and (D) and (E), validate the contribution of the Graph Encoder (GE).
+Additionally, comparisons between (B) and (D), and (C) and (E), highlight the impact of MSVR.
 Ablation of $\beta$.
-Table 5 summarizes the ablation study on the weighting coefficient $\beta$ of the proposed loss function.
-The results indicate that setting $\beta=0.3$ yields the best performance for the current model and dataset.
-Additional ablation studies are provided in the supplementary material.
+Table 5 summarizes the ablation study on the weighting coefficient $\beta$ in the proposed loss function.
+The results show that $\beta=0.3$ yields the best performance for the current model and dataset.
 # Settings Problem 1 Problem 2 Problem 3
@@
 (E) $\checkmark$ $\checkmark$ $\checkmark$ 0.607 1.144 0.598
-Table 4: Ablation study of proposed modules, showing results for EPE ($\downarrow$). $\checkmark$ indicates module presence, $\times$ indicates absence. Setting (G) is our final model.
+Table 4: Ablation study of proposed modules, showing results for EPE ($\downarrow$). $\checkmark$ indicates module presence, $\times$ indicates absence. Setting (E) is our final model.
 # $\beta$ Problem 1 Problem 2 Problem 3
@@
 Conclusion
-We explore a novel approach to fluid motion estimation using spike cameras and propose a specialized network architecture, termed Spike Imaging Velocimetry (SIV), designed to solve PIV tasks while adapting to the scenario-specific characteristics of fluid scenes.
-We propose a Detail-Preserving Hierarchical Transform module to preserve particle signal information during the pre-processing of spike streams.
-We design a Graph Encoder to extract contextual features from highly unstructured and turbulent flow fields.
-We develop a Multi-scale Velocity Refinement module to refine the velocity field.
-We also introduce a challenging and realistic PIV dataset, named Particle Scenes with Spike and Displacement (PSSD).
-Experimental results demonstrate that our method achieves state-of-the-art performance on the PSSD dataset and accurately estimates velocities in small-scale yet critical flow structures such as shear layers and vortices.
-In summary, this study demonstrates the potential of spike cameras in fluid velocity measurement and offers new perspectives and directions for PIV network design.
+We propose a novel approach for fluid motion estimation using spike cameras and introduce a specialized network architecture, Spike Imaging Velocimetry (SIV), designed to address PIV tasks while adapting to the specific characteristics of fluid scenes.
+We develop the Detail-Preserving Hierarchical Transform (DPHT) module to preserve particle signal information during spike stream preprocessing, a Graph Encoder (GE) to extract contextual features from highly unstructured and turbulent flows, and a Multi-scale Velocity Refinement (MSVR) module to refine the velocity field.
+We also introduce the challenging and realistic Particle Scenes with Spike and Displacement (PSSD) dataset.
+Experimental results demonstrate that SIV achieves state-of-the-art performance on the PSSD dataset, accurately estimating velocities in small-scale yet critical flow structures such as shear layers and vortices.
+In summary, this study highlights the potential of spike cameras for fluid velocity measurement and offers new perspectives and directions for PIV network design.
+This paper aims to demonstrate that spike cameras are a promising tool for PIV estimation.
+However, given the complexity of fluid flows, there remains substantial room for improvement: (i) although we have generated a large dataset, its scale and diversity should be further increased; (ii) noise should be modeled more comprehensively to achieve stronger performance on experimental data; and (iii) improving inference efficiency to enable real-time prediction would benefit both laboratory and industrial applications.
+Acknowledgements
+This work is supported by the Postgraduate Research & Practice Innovation Program of Jiangsu Province (No.
+KYCX25_0173), the National Natural Science Foundation of China under Grants 62371006, 62302016, U24B20140, the Jiangsu Association for Science and Technology (JAST) Young Elite Scientists Sponsorship Program (No.
+JSTJ-2025-730), and the Beijing Natural Science Foundation (No.
+3242008).